### v1

# StoneDetector: Conventional and versatile code clone detection for Java

Abstract
Copy & paste is a widespread practice when developing software and, thus, duplicated and subsequently modified code occurs frequently in software projects. Since such code clones, i.e., identical or similar fragments of code, can bloat software projects and cause issues like bug or vulnerability propagation, their identification is of importance. In this paper, we present the StoneDetector platform and its underlying method for finding code clones in Java source and Bytecode. StoneDetector implements a conventional clone detection approach based upon the textual comparison of paths derived from the code’s representation by dominator trees. In this way, the tool does not only find exact and syntactically similar near-miss code clones, but also code clones that are harder to detect due to their larger variety in the syntax. We demonstrate StoneDetector’s versatility as a conventional clone detection platform and analyze its various available configuration parameters, including the usage of different string metrics, hashing algorithms, etc. In our exhaustive evaluation with other conventional clone detectors on several state-of-the-art benchmarks, we can show StoneDetector’s performance and scalability in finding code clones in both, Java source and Bytecode.
[…]
1 Introduction
[…]
int func(int[] array) {
int sum = 0;
for (int i = 0; i<array.length; i++)
if (array[i]>0)
sum *= array[i];
return sum;
(a) Original code fragment
int func( int[] array ) {
int sum = 0;
for ( int i = 0; i < array.length; i++ )
if ( array[i] > 0 )
sum *= array[i];
return sum;
(b) Type 1 clone
int func(int[] arr) {
int s = 0;
for (int i = 0; i<arr.length; i++)
if (arr[i]>0)
s *= arr[i];
return s;
(c) Type 2 clone
int func(int[] arr) {
int s = 0;
for (int i = 0; i<arr.length; i++)
if (arr[i]>0)
s *= (arr[i]*fu());
return s;
(d) Type 3 clone
int func(int[] array) {
int sum = 0;
int i = 0;
while (i<array.length) {
sum = (array[i]>0)?sum*array[i]:sum;
i = i+1;
return sum;
(e) Type 4 clone
Figure 1: Variations of code clones: Examples of Type 1, 2, 3, and 4 code clones
Code clones [1] are identical or similar fragments of code that, e.g., arise due to copy-and-paste practices in software development. Since such practices are quite common, code clones can be frequently found, as empirical studies on code duplication have shown in the past [2, 3]. While code cloning may initially speed up software development or allow for reusing working code, it also introduces long-term challenges: Duplicated code inflates code bases and increases the maintenance burden, as developers have to manually track and update all copies for fixing bugs or implementing code changes. The effects of code cloning are thus controversial [4, 5] and detecting code clones is a vital step toward maintaining code quality, reducing technical debt, and improving program comprehension.
Code clones can come in different variations, ranging from exact replicas of code, over code segments that have been modified in structure or syntax, to fragments that have a completely different syntax but share the same computations (see also Fig. 1). In the literature, code clones are typically categorized into four types based on their similarity [1]: Type 1 clones, which are exact duplicates of code fragments that only differ in layout, comments, or formatting (cf. Fig. 1 (b)). Type 2 clones in addition to Type 1 clones include variations such as identifier renaming or literal changes (cf. Fig. 1 (c)). Type 3 clones go another step further and allow for some inserted, altered, or deleted statements (cf. Fig. 1 (d)). Finally, Type 4 clones are syntactically distinct but perform the same or similar computations (cf. Fig. 1 (e)). Code clone detection has been actively researched [6, 7, 1, 8] and the state-of-the-art approaches for clone detection perform quite well at finding exact and syntactically very similar code clones, i.e., clones of Type 1-2 and part of Type 3 [9]. In contrast, detecting the harder-to-find clone variations of Type 3 and Type 4, e.g., large-gap clones, which include many syntactical modifications [10], or semantic clones, which feature low syntactical similarity but share the same computations [11], is an open and challenging research problem. In this paper, we address this problem and are interested in advancing the detection of code clones with larger syntactical variance.
[…]
1private static void zipDir(File zipDir, ZipOutputStream zos, String path) {
2 final String [] dirList = zipDir.list();
3 final byte [] readBuffer = new byte[chunk];
4 int bytesIn = 0;
5 for (int i = 0; i < dirList.length; i ++) {
6 final File file = new File(zipDir, dirList[i]);
7 final String filePath =
8 ((path != null && path.length() > 0) ? (path+File.separator) : ””)
9 + file.getName();
10 if (file.isDirectory()) {
11 zipDir(file, zos, filePath);
12 continue;
13 }
14 final FileInputStream fis = new FileInputStream(file);
15 try {
16 final ZipEntry anEntry = new ZipEntry(filePath);
17 zos.putNextEntry(anEntry);
18 while ((bytesIn=fis.read(readBuffer)) != - 1) {
19 zos.write(readBuffer, 0, bytesIn);
20 }
21 zos.flush();
22 } finally {
23 fis.close();
24 }
25 }
26}
[…]
1public static void zipDir(File zipDir, ZipOutputStream zos, String path) {
2 String [] dirList = zipDir.list();
3 byte [] readBuffer = new byte[2156];
4 int bytesIn = 0;
5 for (int i = 0; i < dirList.length; i ++) {
6 File f = new File(zipDir, dirList[i]);
7 if (f.isDirectory()) {
8 zipDir(f, zos, path + f.getName() + ”/”);
9 } else {
10 FileInputStream fis = new FileInputStream(f);
11 ZipEntry anEntry = new ZipEntry(path + f.getName());
12 zos.putNextEntry(anEntry);
13 while ((bytesIn=fis.read(readBuffer)) != - 1) {
14 zos.write(readBuffer, 0, bytesIn);
15 }
16 fis.close();
17 }
18 }
19}
Figure 2: Example of two code fragments that are more difficult to identify as code clone
An example of such a harder-to-detect code clone is shown in Fig. 2, taken from the state-of-the-art clone detection benchmark BigCloneBench [12, 13]. Conventional clone detectors have difficulties in identifying the code clone due to the larger syntactical difference, e.g., NiCad [9, 14] is only able to determine the code clone accepting up to 58% syntactical difference, i.e., threshold value 0.58, which is considered to constitute a Type 4 code clone according to the literature [12]. In contrast, our proposed method and clone detection tool, StoneDetector, finds the code clone with a similarity threshold of just 0.26. A closer look at the example reveals that the larger deviation between both code fragments is to a large extent due to different syntactical constructs implementing similar control flow. For instance, the continue in line 12 of the upper code fragment in Fig. 2 is replaced by the else branch in the lower code fragment and the try-finally in lines 15-24 of the upper fragment is replaced by an instruction sequence with the finally block as the sequence’s last instruction in the lower fragment of Fig. 2. In order to find such harder-to-detect code clones, we present in this paper a method for clone detection working on dominator trees. This method exploits the semantics of the code fragments’ control flow for a more permissive textual comparison by encoding and comparing paths of the code fragments’ dominator trees. Dominator trees are a traditional code representation, used in compilers or program analysis, that provides us with a more informed code abstraction when compared to the raw code, token sequences, or syntax trees as used in other clone detectors. As a consequence, the method is able to provide better performance in the detection of code clones with larger syntactical variance, i.e., Type 3 and Type 4 code clones.
The presented clone detection method is implemented in the StoneDetector platform, which is publicly available [^1] https://stonedetector.fmi.uni-jena.de/ and allows for finding code clones in Java source and Bytecode. Using the platform enabled us to perform an in-depth evaluation of the method’s performance on various state-of-the-art benchmarks, including BigCloneBench[12, 13], Google Code Jam, Project CodeNet [15], and GPTCloneBench [16]. As a result, we found that our clone detection method and its implementation in the StoneDetector platform allow for finding exact and near-miss code clones in Java source code at competitive performance with other conventional code clone detectors, including NiCad [14], iClones [17], SourcererCC [18], Deckard [19], CCAligner [10], NIL [20], and Oreo [11]. Additionally, StoneDetector in particular excels at finding source code clones with larger syntactical variance at high precision. StoneDetector also scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art. Furthermore, we have experimented with different configuration parameters for clone detection, e.g., threshold value, code clone size, applied string metrics, and hashing algorithms, etc., and thereby show the versatility of the StoneDetector platform besides identifying its optimal configuration. Eventually, we analyzed the method’s potential in also finding code clones in compilation artifacts like Java Bytecode, in cases where the source code is unavailable in the first place. In summary, the contributions of this paper are:
  • 1.
    We demonstrate the feasibility of finding code clones in Java code fragments by encoding and textually comparing paths of the code fragments’ dominator trees.
  • 2.
    We provide the publicly available StoneDetector platform with an implementation of our code clone detection method as a versatile conventional code clone detector and evaluate its performance and scalability.
  • 3.
    In its evaluation, we can validate that StoneDetector finds exact and near-miss Java source code clones at competitive recall and precision and is able to find even more of the harder-to-find clones with larger syntactical variance, when compared to other conventional code clone detectors on state-of-the-art benchmarks.
  • 4.
    We provide an analysis of the usage of different string metrics for clone detection within StoneDetector, considering string metrics like Levenshtein, Needleman-Wunsch, Hamming distance, and longest common subsequence (LCS)-based metric.
  • 5.
    We describe a variant of StoneDetector, which in particular performs well on large-gap code clones by using a combination of a modified LCS metric with locality sensitive hashing (LSH).
  • 6.
    We study the role of various additional configuration parameters, e.g., threshold value, minimal size of code clones, etc., on clone detection with StoneDetector to empirically derive its optimal configuration.
  • 7.
    We provide a variant of StoneDetector that works on Java Bytecode and experiment with its stack- and register-based intermediate representations.
This work is based upon the authors’ prior conference paper on code clone detection presented at the 2021 IEEE International Conference on Software Maintenance and Evolution [21] and extends it in various ways. First, an analysis of the usage of different string metrics within StoneDetector has been added to the paper, including distance metrics like Levenshtein, Needleman-Wunsch, Hamming distance, and longest common subsequence (LCS)-based metric. Among others, the analysis revealed the versatility of StoneDetector as a clone detection framework: For example, using LCS provided the best tradeoff in terms of recall, precision, and scalability. In contrast, using the Hamming distance yielded a very fast clone detector with impaired performance in finding harder-to-detect Type 3/4 code clones. Second, we now derive the optimal configuration of StoneDetector for clone detection by means of an in-depth empirical analysis of its several configuration parameters (threshold value for path comparison, etc.). Furthermore, expanding the approach for detecting large-gap and subclones, as presented in our previous conference paper [21], we added locality-sensitive hashing (LSH) and a modified LCS metric to StoneDetector, which resulted in performance gains for these types of code clones. We now also include a more exhaustive analysis of StoneDetector’s performance in finding harder-to-detect Type 3/4 code clones by including three more state-of-the-art benchmarks for code clone detection besides BigCloneBench, i.e., Google Code Jam, Project CodeNet, and GPTCloneBench, and two more clone detectors, i.e., iClones and NIL, in the evaluation. Moreover, we also extended our previous approach in Amme et al. [21] for clone detection on stack-based and register-based Java Bytecode representations, such that the StoneDetector platform can be used for both finding code clones in Java source code as well as Java Bytecode.
The rest of the paper is structured as follows: Our clone detection approach and its underlying code clone detection method based on encodings of paths in the dominator tree and string metrics for path comparison are described in detail in Sect. 2. Sect. 3 introduces the StoneDetector platform’s architecture and implementation. In Sect. 4, we present the in-depth evaluation of our approach on various benchmarks for Java code clone detection, including the comparison with state-of-the-art clone detectors and the thorough analysis of the StoneDetector platform’s various configuration parameters. Related work on code clone detection is discussed in Sect. 5 and Sect. 6 finally concludes the paper.
2 Clone Detection using Dominator Trees
The clone detection method described in this paper performs textual comparisons of code fragments’ syntax similar to other text-based or lexical approaches for code clone detection, though the method additionally employs comparisons of the fragments’ control flow at a higher level of abstraction. Moreover, we do not use abstract syntax trees or control flow graphs for this purpose, as in other tree- or graph-based approaches, but the more specialized code representation of dominator trees. Dominator trees [22] are usually utilized in compilers to represent the execution order of instructions. They feature the advantage of, on the one hand, modeling control flow on a higher level when compared to plain syntax or abstract syntax trees. On the other hand, dominator trees support straightforward linearization due to the fact that each node is reachable via one unambiguous path and due to the absence of cycles, in contrast to control flow or program dependence graphs.
2.1 Control flow and Dominator Trees
A control flow graph $G$ is a directed graph defined as $G=(N,E,s)$ with $(N,E)$ denoting the graph and $s\in N$ the start node [22]. We encode the instructions of a code fragment using the graph’s nodes. Note that we here use a single node for each instruction instead of using basic blocks. An edge $(k,l)\in E$ in a control flow graph connects two nodes $k$ and $l$ iff the instruction denoted by $l$ can be immediately executed after the instruction denoted by $k$. Each node $n\in N$ is then reachable via a path from the start node $s$, i.e., a sequence of consecutive edges $e\in E$ that begin at $s$ and end at $n$. An example of a control flow graph representing the lower code fragment shown in Fig. 2, i.e., function zipDir, can be found in Fig. LABEL:fig2 (a).
A dominator tree [22] is a tree and models a certain control flow property of code fragments, specifically, which instruction precedes the other under all possible executions of the respective code fragment. Thus, let the triple $G=(N,E,s)$ denote a certain control flow graph. A node $l\in N$ dominates a node $k\in N$ iff $l$ occurs on all paths from the start node $s$ to node $k$. Accordingly, every node $n\in N$ dominates itself by definition. A node $l\in N$ furthermore strictly dominates another node $k\in N$ iff $l$ dominates $k$ and $l\neq k$. Node $l$ is also immediately dominating $k$ iff $l$ strictly dominates $k$ and there exists no node $m\in N$, such that $m$ strictly dominates $k$ and $l$ strictly dominates $m$. The dominator tree for the control flow graph $G$ is then a directed graph $D=(N,E^{*})$ with $E^{*}=\{(l,k)\,|\,l\mbox{ immediately dominates }k\}$. An example of a dominator tree is visualized in Fig. LABEL:fig2 (b). As can be seen, the shown dominator tree corresponds to the control flow graph in Fig. LABEL:fig2(a). Note that in the dominator tree, each node is reachable from the start node, i.e., the tree’s root, via a single unambiguous path. This is in contrast to the given control flow graph. Also note that there are no backward edges encoding loop edges, and therefore cycles, in the dominator tree, since the loop body cannot dominate the loop condition according to the above definition.
We use an abstract, text-oriented encoding of instructions for our clone detection method, yielding the abstract dominator tree which is depicted in Fig. LABEL:fig2 (c). Therein, instructions and their respective operands are denoted in preorder notation. Furthermore, the distinct variable and field names, as well as literals, are not distinguished and mapped to a single joint symbol, respectively. Thus, all variable names are subsumed by symbol $V$, literals by symbol $L$, and so on, much like in the clone detector NiCad [9]. For example, consider instruction i = 0 which is encoded by [=,V,L] in the abstract dominator tree (cf. Fig. LABEL:fig2). Function calls are optionally represented using a constant pool. A constant pool access is denoted by the character # and the respective function’s entry in the constant pool, as, e.g., in [CALL,#6,V] for instruction zos.putNextEntry(anEntry) (again, cf. Fig. LABEL:fig2). By using abstract dominator trees, we are able to more easily match code fragments that do not completely align with each other syntactically, in particular in the case of Type 2 and part of Type 3/4 code clone deviations.
2.2 Finding Similar Code Fragments
Finding matching code fragments by using graph matching algorithms on their respective dominator trees implies expensive computational costs and is rather inefficient [23]. Instead of using a tree- or graph-based comparison of dominator trees, we have chosen a text-based approach. Accordingly, we linearize the dominator trees into sets of node sequences and, to this end, derive the set of paths starting at a dominator tree’s root node, which are subsequently encoded into sequences and used for the comparison.
Our approach is similar to the method proposed by Alon et al. [24], where they use a linearization of abstract syntax trees based on paths that start and end at distinct leaf nodes of an abstract syntax tree. Though, in contrast, in our method, given a dominator tree $D=(N,E^{*})$, we consider all paths starting at the tree’s root node $s\in N$ and ending at a respective leaf node $l\in\{n\in N\,|\,\not\exists(n,m)\in E^{*}\}$. The resulting set of these paths can then be used to unambiguously represent the dominator tree $D$. A single path is encoded by simply concatenating the (abstract) instructions denoted by the path’s nodes into a sequence of instructions. For example, considering the abstract dominator tree in Fig. LABEL:fig2 (c), the path from the start node to the rightmost leaf node is encoded by the sequence:
[…]
Let $D$ be a dominator tree, $\mathit{paths}_{D}$ the set of its paths, and $\mathit{leaves}_{D}$ the set of its leaf nodes in the following. The set of paths unambiguously representing $D$ is given by: $D_{\mathit{set}}=\{s\rightarrow\cdots\rightarrow n_{k}\,|\,s\cdots n_{k}\in\mathit{paths}_{D},n_{k}\in\mathit{leaves}_{D}\}$. We call the set $D_{\mathit{set}}$ the description set of dominator tree $D$. In the case of the abstract dominator tree depicted in Fig. LABEL:fig2 (c), we have four paths starting at the start node and ending in the leftmost, the left inner, the right inner, and the rightmost leaf node, respectively. Accordingly, the description set shown in Fig. LABEL:fig2 (d) contains the four instruction sequences representing the paths.
Given two dominator trees $D$ and $D^{\prime}$, which include two code fragments, $f$ and $g$, respectively, we call $f$ and $g$ strictly similar iff both description sets $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$ have the same number of elements and there exists a similar path $p^{\prime}\in D^{\prime}_{\mathit{set}}$ for each path $p\in D_{\mathit{set}}$. In cases where only the latter condition is satisfied, meaning that $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$ do not share the same number of elements, $f$ and $g$ are called partially similar.
Note that we have not defined the similarity of paths, though. The similarity of two paths included in the dominator trees $D$ and $D^{\prime}$, respectively, can be formally assessed by a distance function $\delta:paths_{\mathit{D}}\times paths_{\mathit{D^{\prime}}}\rightarrow\mathbb{N}$ determining the distance of two paths. Basically, the distance function $\delta$ can be freely chosen according to its requirements. The function only has to satisfy the metric conditions, i.e., returning zero in case of identical paths, a value greater than zero for distinct paths, being symmetrical in its arguments, and fulfilling the triangle inequality. The distance function is later used as a metric for path similarity of dominators tree and subsequently also for assessing the similarity of code fragments and finding code clones.
2.3 Metrics for Path Comparison
For a naïve comparison of two paths $p$ and $q$ of dominator trees, we could simply analyze their respective nodes, i.e., instructions, for textual equality in the order of their occurrence. Such a comparison, though, assumes that only those paths are similar that share the same nodes; in other words, paths in which contained instructions are textually identical. Thus, a more permissive approach is required, in particular in the case of Type 3 and Type 4 code clones, where the paths of the dominator trees may deviate from one another in some of their respective nodes. Even more, a path may be embedded in the other matching path in case of large-gap or subclones [21]
A more permissive approach can be implemented using string metrics, which are used in approximative string matching and are therefore excellent candidates for our above-defined distance function $\delta$. Since string metrics are based on text strings, we depict the nodes contained in a path as individual characters, and two nodes are then considered the same character if they represent the same instruction. As a result, we can represent a single path of an (abstract) dominator tree as a text string by concatenating the respective characters of the path’s nodes in the order of their occurrence. String metrics are then readily applicable for the path comparison.
The Hamming distance [25] is a basic string metric, which is defined on text strings with fixed length. This distance measures the minimum number of substitutions to change one string into another and thus yields the number of deviating characters in two given strings. Therefore, as a limitation, the text strings are only compared up to the length of the respective shorter string, and the number of remaining characters in the longer string is merely added to the so far calculated distance measure. Note that the conventional Hamming distance therefore tends to penalize paths with different lengths.
Example:
  • Code fragment 1: for (int j=0; j<<m; j++) { acc=acc-j; proc(acc); }
  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
  • Code fragment 2: for (int l=0; l<<m; l++) { acc=acc+l; }
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow$ [POSTINC,V]
Hamming:
  • 1.
    Substitute [=,V,+,V,V] with [=,V,-,V,V]
    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[POSTINC,V]}$
  • 2.
    Substitute [POSTINC,V] with [CALL,#4,V]
    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$
  • Hamming distance $\delta_{\mathit{H}}(p,q)=3=2+1$ because $p$ has one more character than $q$.
Levenshtein:
  • 1.
    Substitute [=,V,+,V,V] with [=,V,-,V,V]
    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[POSTINC,V]}$
  • 2.
    Insert [CALL,#4,V]
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[\mbox{CALL},\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
  • Levenshtein distance $\delta_{\mathit{L}}(p,q)=2$.
Needleman-Wunsch:
  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$ $\rightarrow\texttt{[POSTINC,V]}$
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[POSTINC,V]}$
  • Needleman-Wunsch-based distance $\delta_{\mathit{NW}}(p,q)=max(|p|-score(p,q),|q|-score(p,q))=2$, where $score(p,q)=3$.
Longest Common Subsequence (LCS):
  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\ \rightarrow\texttt{[POSTINC,V]}$
  • $z=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[POSTINC,V]}$
  • LCS-based distance $\delta_{\mathit{LCS}}(f,g)=max(|p|-|z|,|q|-|z|)=2$.
Figure 4: Distance metrics for calculating path similarity
A number of string metrics can be borrowed from bioinformatics, where they are usually used for DNA sequence alignment. One such metric is the Levenshtein distance [26]. This distance measure is defined as the minimum number of mutating operations for two given text strings, i.e., inserting, deleting, or substituting a single character, which are needed to convert the one string into the other. In other words, the Levenshtein distance is an extension to the previously described Hamming distance, where text strings are not limited to having the same length, and the insertion and deletion operations are additionally allowed besides character substitution.
Another metric that has its origins in the alignment of DNA sequences is the Needleman-Wunsch algorithm [27]. The algorithm computes an alignment of two given text strings that is optimal with respect to a certain scoring scheme using two steps. In the first step, a similarity matrix of the two text strings is generated based upon the scoring scheme, which defines weights for character matches, mismatches, and gaps. In the second step, the optimal string alignment with maximal total score is computed by performing a backward traversal on this similarity matrix.
The application of string metrics to the comparison of paths in abstract dominator trees is illustrated in Fig. 4 by way of an example. In the upper part, two code fragments are shown, conjoined with two paths $p$ and $q$ from their abstract dominator trees. As can be seen below, the (abstract) instructions take on the role of characters in the visualized calculations of string metrics. In the alignment provided for the Needleman-Wunsch algorithm, vertical lines between nodes represent matches and horizontal lines represent inserted gaps. For the value of $store(p,q)$, a scheme is used in which the weights for matches are set to $1$, for mismatches to 1-1, and for gaps to 0.
The Longest Common Subsequence (LCS) algorithm [28, 9] is a well-known tool for comparing the contents of files, used for example in the Unix diff utility or in the revision control system git. Besides, clone detectors like NiCad are relying on the LCS algorithm as well. In our case, the algorithm is applied to paths instead of lines of code, such that the longest common subpath is computed, e.g., $\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[POSTINC,V]}$ for paths $p$ and $q$ in Fig. 4. This longest common subpath then forms the basis for defining the metric $\delta$ for two paths of dominator trees, which equals the maximum number of unique nodes for both paths.
Using the original LCS-based metric, though, can become problematic in cases of large-gap code clones [10], e.g., code clones where one code fragment is embedded into the other. The LCS metric then has issues in recognizing the similarity of the respective paths of the code fragments’ dominator trees, since the paths differ too much in their length. For this reason, a modified variant of the LCS-based metric can be used, which performs additional similarity checking in situations where two paths are not considered similar by the original LCS-based metric and analyzes whether or not the derived LCS distance value is close to the length of a respective smaller path. On a technical note, code fragments consisting of only one basic block and thus path are then considered as code clones if the calculated LCS value corresponds to a value lower than 30% of the number of nodes of the smaller code fragment. Furthermore, a path $p$ is also considered a subpath of a path $q$ if the calculated LCS value of $p$ and $q$ is less than 10% of the length of $p$. Finally, the smaller of both paths is required to have a minimum length of 10 nodes.
Note that there exist other string metrics as well, besides the ones discussed above, which have also been implemented in the StoneDetector platform. Among these are, e.g., N-gram distance, cosine distance, Jaccard similarity. For the sake of brevity, we here limit ourselves to the subset of metrics that we use in the evaluation in Sect. 4. For a more comprehensive discussion of string metrics, we refer, e.g., to the survey [29].
2.4 Fingerprints and Hashing
However, using string metrics on the plain paths of (abstract) dominator trees for path comparison is rather inefficient due to the need for node matching, i.e., matching equal instructions, multiple times for a single comparison of two paths. Therefore, to minimize the costs, we conduct the node matching on fingerprints instead, which encode the respective nodes’ instructions. Such a fingerprint is simply the hashed value of an instruction, for which the string metrics are adjusted accordingly. There are apparently various hashing algorithms available for this purpose. For example, the MD5 algorithm [30] provides a widely used hash function with 16-byte hash values, which is often used as a checksum of file downloads and suffices for creating collision-free encodings of instructions in our abstract dominator trees. The MD5 algorithm is the default hashing algorithm for fingerprints in the StoneDetector platform, and its application for encoding the description sets of the example in Fig. LABEL:fig2 is shown in the lower part of the same figure.
We have implemented and experimented with two other hashing algorithms as well, including a simple 4-byte hash function based upon prime number encoding and Locality-Sensitive Hashing. For the latter, we observe that the MD5 hashing algorithm deliberately generates different hash values for distinct instructions in the description sets in order to avoid collisions. For instance, the two instructions i=-i*2 and i=i*-2 are mapped to different hash values even though they are syntactically and semantically very similar statements. As a result, similar paths and description sets might not be identified as such, and code clones might therefore be missed.
By contrast, Locality-Sensitive Hashing (LSH) [31] can be used to create identical hash values for similar elements. LSH places elements to be mapped to hash values into buckets so that similar elements are most likely to be stored in the same bucket. Within the algorithm, the number of buckets is always chosen so that it is significantly smaller compared to the number of elements. In previous work, LSH has already been used for code clone detection (cf. Alomari and Stephan [32]). For that purpose, the number of comparisons between code fragments is reduced by placing similar fragments in the same bucket. When similarity of code fragments is later analyzed, only those fragments in the same function have to be considered.
In StoneDetector, LSH can be used as an alternative to MD5 as well, though at a more fine-grained level. Using LSH allows for assigning the same hash values to similar nodes. To this end, the nodes contained in description sets are clustered according to the length of the represented instruction, and the similarity-preserving hash values are determined separately for the resulting clusters. By default, StoneDetector uses the length clusters 3 to 8. Instructions of the corresponding length are assigned to the length clusters, whereby instructions with a length greater than 8 are also assigned to the length cluster 8. LSH generates 200 buckets for each of the different length clusters, in which the instructions assigned to the individual length clusters are then organized according to their similarity.
2.5 Comparison of Description Sets
Choosing one of the metrics in combination with a hashing algorithm, as presented in Sect. 2.3 and Sect. 2.4, for the distance function $\delta$ allows us to compare description sets and subsequently assess the similarity of code fragments. To this end, given two code fragments with respective description sets $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$, the most similar path $q$ of $D^{\prime}_{\mathit{set}}$ is identified for each path $p$ of $D_{\mathit{set}}$, such that $q$ has a minimal value $\delta(p,q)$, and the arithmetic mean for these values over all paths $p\in D_{\mathit{set}}$ is computed. Formally, we can define the distance of the two description sets $D_{\mathit{set}},D^{\prime}_{\mathit{set}}\in\mathbb{D_{\mathit{set}}}$, having $|D_{\mathit{set}}|\leq|D^{\prime}_{\mathit{set}}|$ without loss of generality, as a function $\Delta$ such that:
$\Delta:\mathbb{D_{\mathit{set}}}\times\mathbb{D_{\mathit{set}}}\rightarrow\mathbb{N}$, with
$\Delta(D_{\mathit{set}},D^{\prime}_{\mathit{set}})=\sum\limits_{p\in D_{\mathit{set}}}min(\{\delta(p,q)\,|\,q\in D^{\prime}_{\mathit{set}}\})/|D_{\mathit{set}}|$
We are, though, not interested in whether two description sets are identical, rather, we want to find code fragments having similar description sets, i.e., code clones. A threshold $\tau$ is therefore defined, limiting the maximum difference of description sets that are considered as similar. Using $\tau$, we have $\Delta(D_{\mathit{set}},D^{\prime}_{\mathit{set}})<\tau$ for two description sets $D_{\mathit{set}},D^{\prime}_{\mathit{set}}\in\mathbb{D_{\mathit{set}}}$ iff $D_{\mathit{set}},D^{\prime}_{\mathit{set}}$ are similar and their corresponding code fragments constitute a code clone.
3 Architecture and Implementation
The approach to code clone detection introduced in Sect. 2 has been implemented in the StoneDetector platform [21, 33] thus providing a full-scale stand-alone clone detector for Java. As input, StoneDetector expects a number of Java source or Bytecode files for which the platform generates a list of all code clone pairs, which are identified and mapped back to the original code by file location and line numbers. The resulting information can then be used, for example, to search for similar programs in code repositories [34], to analyze the evolution of code [35], or to track bugs and vulnerabilities propagated via duplicated code [36].
Figure 5: Overview of the StoneDetector platform
An overview of the StoneDetector platform can be seen in Fig. 5, which shows its inner structure and how its components interact with each other. Considering, to begin with, Java source code, the platform starts with performing lexical and syntax analysis on the given Java source files. This step yields the abstract syntax trees, which are subsequently subject to control flow analysis for generating (abstract) dominator trees. Afterward, paths are extracted from the generated dominator trees and linearized into description sets. In case fingerprints are enabled, the individual nodes of the description sets are further encoded using a hashing algorithm, which yields fingerprint sets. Either raw description sets or fingerprint sets are eventually transferred to the clone matcher, which checks for similarities using one out of the several available string metrics. Based upon this similarity analysis, code clones are identified and returned for further processing. We will next provide further details on the individual components of the described processing pipeline for Java source code files:
Lexical/Syntax Analysis
The lexical and syntax analysis in the Java source code frontend of the StoneDetector platform is utilizing the Spoon analysis framework [37] [^2] https://spoon.gforge.inria.fr/. The Spoon framework includes libraries for parsing Java source code files up to Java version 20 into abstract syntax trees (AST) and provides the AST metamodel and functionality for manipulating them.
Control Flow/Dominator Analysis
The Spoon framework also supports generating control flow graphs of individual Java functions based on its AST model. We can therefore resort to Spoon’s control flow builder for our control flow analysis. However, the framework currently includes neither a meta-model for dominator trees nor any dominator analysis. For the construction of dominator trees, the respective builder from the WALA framework [38] [^3] https://github.com/wala/WALA has been integrated into the StoneDetector platform. As before, dominator trees are built at the function level. In addition, due to the coarse abstraction for Java’s exception handling in control flow graphs generated by Spoon, we have implemented a standard interprocedural exception analysis for Java as described by Chang et al. [39]. Using the analysis allows for a more fine-grained control flow mapping of Java’s exception handling. The analysis is, though, optional, but experiments have shown that enabling exception analysis implies better results for clone detection.
Path Extraction and Description Set Encoding
As described in Sect. 2, dominator trees are transformed into description sets in order to subsequently perform textual clone matching. To this end, a (abstract) dominator tree is traversed such that the set of its paths starting at the root node and ending at leaf nodes, respectively, is generated (cf. Fig. LABEL:fig2). Note that we in general perform identifier and literal abstraction in the resulting description sets, much like as in other clone detectors, e.g., NiCad [9]. Though function names at call sites may optionally be conserved, which provided better experimental results in some cases. We therefore offer a switch for enabling/disabling identifier abstraction for function names at call sites. Likewise, the StoneDetector platform treats encoding of description sets into fingerprint sets via a hashing function as optional. Raw description sets can thus be directly committed to the clone matcher or beforehand transformed such that the contained instructions are encoded using one of the several implemented hashing algorithms, i.e., MD5, 4-byte prime number hashing, or LSH.
Clone Matching
The clone matching component receives a set of description or fingerprint sets as input and performs pairwise similarity checks. As described in Sect. 2, the similarity of description sets, i.e., code fragments, is traced back to the accumulated similarity of its paths in our approach. Path similarity is computed using one out of several string metrics available in the StoneDetector platform. Among the currently available metrics are the Hamming distance [25], both the normal and the weighted Levenshtein distance [26], the Needleman-Wunsch-based metric [27], as well as the original or modified LCS-based metric [28]. Further supported metrics include the N-Gram, Jaccard, cosine, Jaco-Winkler, and Damerau-Levenshtein metrics.
The thus defined processing pipeline in the StoneDetector platform is parallelized. On the one hand, generation of description sets can be done in parallel with the help of streams. If all description sets have been eventually generated, clone matching begins. This process, on the other hand, can also be done in parallel, again using streams. The configuration of a specific run of the StoneDetector platform allows for the specification of the number of threads to be used in both computations. In order to further increase throughput, especially when processing larger numbers of code fragments, a copy strategy can optionally be used for description set handling. NiCad uses a similar technique for improving its runtime behavior [9, 14]. In StoneDetector, however, the technique has been adopted in such a way that the number of identified clones remains unchanged whether or not the technique is enabled. The basic idea of the copy strategy is to continue with further comparisons only for one of the two code fragments $f$ and $g$ whenever a clone pair $f$ and $g$ is found: Suppose comparisons are continued for fragment $f$ and terminated for fragment $g$. After all comparisons are finished, code clones found for $f$ are added to the clones already found for $g$, up to termination. Note that in contrast to NiCad’s implementation, in our variant of the copy strategy, two code fragments are only considered as a clone pair if they are identical to each other, i.e., if they have a distance value of zero. Using our copy strategy, runtime behavior does not improve as much as in NiCad, however, accuracy of the resulting code clones is guaranteed.
Apparently, as can be seen by the above discussion, the StoneDetector platform provides a rich set of configuration parameters that can be used for optimizing clone detection and demonstrates the platform’s versatility. Among the most critical parameters are the threshold value $\tau$ defining the required similarity of code clones, the minimum size $l$ of code fragments acceptable as code clones, the string metric $m$ used for path matching, and the hash function h used for description set encoding (cf. also Table 1). For an in-depth description of the platform’s configuration parameters, we refer the reader to the supplementary information available online [^4] https://stonedetector.fmi.uni-jena.de/.
Configuration parameter Value range Default value
Similarity threshold $\tau$ $0\leq\tau\leq 1.0$ $0.3$
Minimum clone size in code lines $l$ $0\leq l$ 15
String metric $\delta$ LCS LCS
LCSmodified
Levenshtein
Needleman-Wunsch
Hamming
N-Gram
Jaccard
Cosinus
Jaco-Winkler
Damerau-Levenshtein
Hashing algorithm $h$ none MD5
4-byte hashing
MD5
LSH
Table 1: Main configuration parameters in the StoneDetector platform
3.1 Extensibility of the StoneDetector platform
The use of fingerprints to represent description sets makes clone detection in the StoneDetector platform fundamentally independent of the selected programming language and possible program representation. In fact, as is also depicted in Fig. 5, StoneDetector can be divided into a frontend and backend, with the former generating description sets of given code fragments and the latter performing the actual clone matching. The strict separation of frontend and backend allows us to facilitate a readily extensible platform, which can be easily extended for processing other programming languages as well as other internal program representations, such as control flow graphs, abstract syntax trees, or even just token sequences of linear code.
Such an extension is only required to map the new programming language or program representation to fingerprints by using a certain hash function. The clone matching can subsequently be performed using the existing backend of the StoneDetector platform. The specific hash function used for construction is irrelevant to the functionality of the backend. In this way, fingerprints form a general and elegant interface for extending our system to a wide variety of programming languages and program representations.
To demonstrate the platform’s versatility and extensibility, we have added the handling of Java Bytecode as another frontend of the StoneDetector platform [40] (cf. Fig. 5). The new frontend gets a set of class files as input and transforms them to fingerprint sets, which are committed to the existing clone matcher of StoneDetector. Specifically, processing class files is implemented via the Soot framework [41], which first transforms the contained Bytecode into the Bytecode-like Baf or into the register-based Jimple representation. Afterward, dominator trees are generated based upon the Baf or Jimple representation and paths contained in the dominator trees are extracted and encoded into fingerprint sets, as required for clone matching.
4 Evaluation
In the evaluation, we are interested in finding answers to a number of research questions, in particular about the performance of our code clone detector compared to the current state of the art. We also present results that we believe to provide general insights into code clone detection for Java and to help other researchers when implementing code clone detectors, besides showing the versatility of the StoneDetector platform. Thus, we address the following research questions:
  • 1.
    Research Question 1: How many code clones does our clone detection method and tool StoneDetector detect, i.e., what is its recall compared to the state of the art?
  • 2.
    Research Question 2: How many of StoneDetector’s reported code clones are real code clones, i.e., what is the tool’s precision compared to the state of the art?
  • 3.
    Research Question 3: Are conventional code clone detectors like StoneDetector in general able to find more difficult-to-identify code clones, i.e., code clones with larger syntactical deviation?
  • 4.
    Research Question 4: Does StoneDetector scale to large codebases containing up to multiple million lines of code?
  • 5.
    Research Question 5: How do the used threshold value and the minimal size of code clone influence StoneDetector’s performance?
  • 6.
    Research Question 6: Which role do different string metrics and hashing methods play in a code clone detector like StoneDetector?
  • 7.
    Research Question 7: Can source code clones also be found in Java Bytecode using clone detectors like StoneDetector?
Tool Configuration parameters
Deckard 70% similarity, 2 token stride
SourcererCC function granularity, 70% similarity
NiCad 6.2 blind renaming, literal abstraction,
70% similarity threshold
iClones informat single, minclone 50, minblock 15
CCAligner 60% threshold, 6 window size,
1 edit distance
StoneDetector 70% similarity threshold ($\tau=0.3$),
metric $\delta_{\mathrm{LCS_{modified}}}$, LSH hashing
Oreo 55% action filter, 60% partition threshold
NIL 5-gram, 10 partitions, 10% filter threshold,
70% threshold
Table 2: Tools’ configuration parameters
For our evaluation, we have mainly chosen BigCloneBench [12, 13] [^5] https://github.com/jeffsvajlenko/BigCloneEval as the state-of-the-art benchmark for Java code clone detection. However, in order to avoid an evaluation skewed to this benchmark and to assess the performance of clone detectors for the more difficult-to-identify code clones of Type 3 and Type 4, we also consider the benchmarks Google Code Jam [^6] https://zibada.guru/gcj/, Project CodeNet [15] [^7] https://github.com/IBM/Project_CodeNet, and GPTCloneBench [16] [^8] https://github.com/srlabUsask/GPTCloneBench.
In the comparison of StoneDetector with other code clone detectors for Java, a set of representative state-of-the-art tools has been chosen, including conventional text-based, token-based, tree-based, and data-driven clone detectors, namely, we use the tools:
Benchmark Type 1 Type 2 Very-Strongly Strongly Moderatly Type 4
Type 3 Type 3 Type 3
BigCloneBench 47,146 4,609 4,163 16,631 83,444 8,219,320
GPTCloneBench 5,431 11,558
Google Code Jam 10,936,986
Project CodeNet 11,212,500
Table 3: Benchmarks (clone numbers vary according to chosen minimal clone length)
In the evaluation experiments, we only consider code fragments with a minimal length of 15 lines in the original source code if not otherwise stated. Considering smaller code fragments in general results in a skewed comparison. For instance, in a Java method with only 6 lines of source code, half of the lines are usually just taken by the method’s annotations, signature, and curly braces. However, for the curious reader, we also provide experiments regarding different minimal sizes of code fragments (cf. Sect. 4.6). The clone detectors’ configurations are otherwise chosen in conformance with the literature; see also Table 2. All experiments have been performed on a Linux workstation with an Intel Core i7-10700K processor and 32 GB RAM.
Note that we provide an online version of StoneDetector, the tool’s source code, a runnable container image, and supplementary material, which we here omit due to space constraints [^16] https://stonedetector.fmi.uni-jena.de/ [33, 42, 43, 44].
4.1 Benchmarks
The benchmark BigCloneBench [12] is the state-of-the-art benchmark for code clone detection in Java source code. It contains 8,379,922 code clone pairs over 55,499 Java source files with approximately 15.4 million lines of code. The benchmark has been assembled from real-world software projects of the IJaDataset 2.0 [45]. To this end, methods in the IJaDataset have been heuristically classified and manually validated according to a set of 43 different functionalities, e.g., bubble sort or database update and rollback. If two methods are then classified as having the same functionality, they are considered a code clone pair. The resulting code clones are examined and classified according to their clone type, i.e., Type 1, Type 2, Type 3, and Type 4 code clones, as introduced in Sect. 1. The latter two are further differentiated into:
  • 1.
    Very-Strongly Type 3 with 90-100% syntactical similarity after normalization
  • 2.
    Strongly Type 3 with 70-90% syntactical similarity after normalization
  • 3.
    Moderately Type 3 with 50-70% syntactical similarity after normalization
  • 4.
    Weakly Type 3/Type 4 with less than 50% syntactical similarity after normalization
Note that the benchmark BigCloneBench is quite skewed towards harder-to-find Type 4 code clones, as can be seen in Table 3.
A benchmark including in particular harder-to-detect code clones has been defined with GPTCloneBench [16]. The authors of the benchmark used the code clones of the SemanticBench [46] benchmark in prompts for generating additional code clones with the GPT-3 large language model. Manually validating the generated code clones afterward resulted in a set of 37,14 code clones, of which 16,989 are clone pairs with Java source code that we can use in our experiments. In spite of the benchmark’s authors intentions, we use the 5,431 near-miss code clones of Type 1 and Type 2 as well as the 11,558 Type 3 and Type 4 code clones in the experiments. Google Code Jam was an international programming contest where contestents were solving programming tasks. The posted solutions are publicly available, comprising 3.5 million code fragments in different programming languages for 716 programming tasks. Two different solutions to the same programming task can apparently be considered to share the same computations and thus to constitute a code clone. The thus defined code clones can vary across the different types of code clones. We consider solutions for the year 2022, and, as only part of the solutions are Java code, we are able thus to use 10,936,986 code clone pairs in our experiments. The last benchmark also originates from programming contests. Project CodeNet [15] assembles 14 million code fragments and more than 500 million lines of code with solutions to 4053 programming tasks, written in 55 programming languages. Different solutions to the same programming task are again considered code clones, however, near-miss code clones, i.e., code fragments sharing above 90% syntactical similarity, are omitted in this benchmark [15]. Around 5% of the code fragments, or 25 million lines of code, in Project CodeNet are Java code, which results in 11,212,500 code clone pairs used in our experiments. Note that all three aforementioned benchmarks, in contrast to BigCloneBench, do not further distinguish between the different clone types.
Deckard SourcererCC NiCad iClones CCAligner StoneDetector NIL Oreo
Recall 18,318 20,827 20,829 20,842 20,831 20,828 20,844 20,844
T 1 87.8% 99.9% 99.9% 99.9% 99.9% 99.9% 100% 100%
Recall 3,317 3,444 3,460 3,453 3,456 3,475 3,460 3,475
T 2 95.4% 99.1% 99.5% 99.3% 99.4% 99.9% 99.5% 99.9%
Recall 2,139 3,209 3,243 3,253 3,218 3,209 3,271 3,283
VST 3 65.0% 97.6% 98.6% 98.9% 97.8% 97.6% 99.5% 99.8%
Recall 1,806 5,784 6,246 6,531 5,943 6,808 6,549 6,584
ST 3 24.4% 78.1% 84.3% 88.2% 80.2% 92.0% 88.5% 88.9%
Recall 2,311 1,462 169 4,051 1,784 6,324 3,499 5,187
MT 3 8.6% 5.5% 0.6% 15.1% 6.6% 23.6% 13.1% 19.4%
Recall 57,063 250 7 6,554 2,133 6,061 3,821 6,821
T 4 1.3% 0.0% 0.0% 0.1% 0.0% 0.1% 0.1% 0.1%
#Clones 13,239,074 804,315 754,022 987,659 1,244,160 1,300,458 1,274,852 1,015,425
Precision 46% 100% 100% 97% 98% 99% 99% 99%
Table 4: Recall and precision of clone detectors on Java source code on BigCloneBench (T 1 - Type 1, T 2 - Type 2, VST 3 - Very-Strongly Type 3, ST 3 - Strongly Type 3, MT 3 - Moderately Type 3, T 4 - Type 4, #Clones - total number of detected clones)
4.2 Research Question 1: Recall
In order to address the first research question about how many code clones in Java source code our clone detection method is able to find, we ran StoneDetector on BigCloneBench via BigCloneEval [13]. For a comparison with the state of the art, we also ran seven other clone detectors on the same benchmark. The resulting numbers of identified known code clones, i.e., recall values, for the tools Deckard, SourcererCC, NiCad, iClones, CCAligner, StoneDetector, NIL, and Oreo are shown in Table 4, as measured by BigCloneEval. Since these numbers reflect only the annotated code clones in the benchmark (cf. also Sect. 4.1), we additionally list the overall number of detected clones. As noted before, we only consider clones with a minimum size of 15 lines of code. Duplicate clones, as reported by some of the tools, have been removed. Recall values are given in absolute and relative numbers.
As can be seen in the table, all the code clone detectors, with the exception of Deckard, perform well on exact and near-miss code clones, i.e., code clones of Type 1, Type 2, and Very-Strongly Type 3. The respective recall values are located in the interval 97-100% and only for Deckard do we observe a worse recall value of 65%. These numbers in general reflect the results of prior research [18]. Thus, we can state that StoneDetector performs competitively with the state of the art for exact and near-miss code clones.
As expected, beginning with code clones of Strongly Type 3, recall values are declining due to the larger syntactical differences of code fragments. We can also observe a difference between near-miss code clone detectors, e.g., NiCad and SourcererCC, and tools addressing harder-to-find clones, i.e., CCAligner, iClones, Oreo, StoneDetector, and NIL. StoneDetector achieves the best recall values for code clones of Strongly Type 3 and Moderately Type 3, but is overtaken by the tools Deckard, iClones, and Oreo for code clones of Type 4. However, the high number of Type 4 clones Deckard is finding has to be taken with care, considering its poor precision as described in the following section and also reported in the literature [18, 10].
When compared with state-of-the-art conventional clone detectors for Java source code, StoneDetector detects code clones with competitive recall and finds more code clones with 30% to 50% syntactical variance.
4.3 Research Question 2: Precision
In addition to a high number of identified code clones, a code clone detector shall also avoid reporting on wrong code clones. Thus, we have to measure the clone detector’s precision, i.e., how many of the reported code clones are actual code clones. Unfortunately, answering the second research question is challenging, since BigCloneBench/BigCloneEval only contains information about a subset of all possible code clones in the benchmark. Due to the benchmark’s construction, only code clones of the considered 43 functionalities are known [12]. We have therefore performed a manual evaluation on random samples of the tool’s reported code clones, as is also usually conducted in the respective research literature [10, 11, 18].
To this end, we randomly selected samples of 400 clone pairs of the overall reported code clones for each clone detector. The clone pairs were blindly, i.e., without revealing the respective tool, given to the three authors of this paper, who acted as judges. Each judge individually assessed the given clone pairs as actual or wrong clone pairs (true/false positive). Ambiguous assessments between the three judges were discussed and resolved afterward.
The resulting precision estimates for the clone detectors are also shown in Table 4. As already mentioned, Deckard achieves subpar precision, i.e., only 46% of Deckard’s reported code clones are assessed as such by the three judges. This observation has already been reported for prior evaluations [10, 11]. The other tools, including StoneDetector, perform at high precision levels between 97-100%, while near-miss clone detectors NiCad and SourcererCC achieve the best precision. Compared to prior evaluations, we find in general better precision numbers. This can, though, be attributed to the chosen minimum size of 15 lines of code clones, compared to the 6 lines used in other work. Increased minimum clone size supports the precision of clone detectors, as has already been reported by Wang et al. [10].
StoneDetector detects code clones in Java source code, similar to other state-of-the-art clone detectors, at very high precision.
We conducted further studies on the characteristics of the code clones found by StoneDetector and the other clone detectors. To this end, we made pairwise comparisons of the detected code clones for StoneDetector and each other tool, respectively, and looked into their common and exclusively identified code clones. On the one hand, we were able to observe a large fraction of commonly identified code clones. This mostly points to the subset of exact and near-miss clones as identified by all clone detectors (cf. also Table 4). On the other hand, we also found a significant portion of code clones with larger syntactical variance, in particular structural code clones and subclones, that are exclusively identified by StoneDetector and not by the other tools. For more details, we refer the reader to our previous study in [21].
4.4 Research Question 3: Generalizability
Our third research question addresses the capability of clone detectors to identify code clones in source code with larger syntactical variability, in particular code clones of Strongly Type 3 and beyond. Besides, we are also interested in analyzing the clone detectors’ performance on benchmarks other than BigCloneBench in order to prevent the evaluation from being biased toward this particular benchmark. We therefore ran the clone detectors SourcererCC, NiCad, CCAligner, StoneDetector, NIL, and Oreo on the three benchmarks GPTCloneBench, Google Code Jam, and Project CodeNet, respectively, and report the observed recall values in Table 5. We omitted Deckard and iClones due to their lower precision compared to the other tools. We again limit clones to those with a minimal size of 15 lines, use the same configurations, and depict recall in absolute and relative numbers.
Benchmark SourcererCC NiCad CCAligner StoneDetector NIL Oreo
GPTCloneBench 4,577 3,646 3,452 4,068 4,681 4,366
26.9% 21.5% 20.3% 23.9% 27.6% 25.7%
Google Code Jam 101,843 54,363 102,129 362,989 212,597 115,879
0.9% 0.5% 0.9% 3.3% 1.9% 1.1%
Project CodeNet 104,790 121,271 145,218 713,942 646,232 537,322
0.9% 1.1% 1.4% 6.4% 5.8% 4.8%
Table 5: Recall of clone detectors on GPTCloneBench, Google Code Jam, Project CodeNet
As shown in the table, recall values are in general at the lower end of the spectrum, which is, however, expected due to the nature of the benchmarks (cf. Sect. 4.1). In particular, the benchmarks Google Code Jam and Project CodeNet contain code clones, which originate from heterologously developed code and therefore usually include semantic code clones with differing syntax but similar computations. Thus, the recall values for all tools do not exceed 10% for both benchmarks. We can as well observe a distinction between the near-miss clone detectors NiCad and SourcererCC and tools addressing more difficult code clones, i.e., CCAligner, NIL, Oreo, and StoneDetector, where the latter ones perform better. While the contrast is not so apparent on the benchmark GPTCloneBench, due to its inclusion of Type 1/2 clones, it becomes obvious for the other two benchmarks. We also observe StoneDetector to perform the best on these two benchmarks.
StoneDetector in particular excels at finding source code clones with larger syntactical variance when compared to other conventional state-of-the-art clone detectors on different benchmarks.
Tool 1 million LOC 10 million LOC 100 million LOC
NiCad 57s 18.49min 2.45h
CCAligner 18s 6.58min -
SourcererCC 4.17min 28.09min 46.52h
iClones 24s 7.26min -
StoneDetector 24s 6.17min 7.13h
Oreo 2.42min 9.54min 7.10h
Table 6: Scalability experiments on IJaDataset 2.0
4.5 Research Question 4: Scalability
Scalability is another important requirement for clone detectors, i.e., such tools need to be able to process large quantities of code in reasonable time. In order to evaluate StoneDetector’s capabilities of analyzing large fractions of code, we use a setup described by Sajnani et al. [18]. To this end, random subsets of files from the IJaDataset 2.0 [45] are generated, comprising 1, 10, and 100 million lines of code, respectively. We ran StoneDetector and other clone detectors on these subsets and report their runtime in Table 6. As can be seen, Deckard has been omitted, since Deckard’s memory consumption exceeds available resources, as has also been noted by Sajnani et al. [18]. Furthermore, we observe that CCAligner as well as iClones are not able to scale to 100 million lines of code (cf. also [10]). Also, contradicting previous results in [18], the clone detector NiCad in its recent version NiCad v6.2 scales to more than 100 million lines of code and also provides the lowest runtime. We also see that StoneDetector keeps up well with the other state-of-the-art tools and analyzes 100 million lines of code in 7 hours and 13 minutes and 10 million lines of code in under 7 minutes. Using a different string metric as in the default configuration allows StoneDetector to run even faster but also partially sacrifices recall and precision (cf. Sect. 4.7).
Due to these encouraging results, we performed another scalability check using StoneDetector on the entire IJaDataset 2.0 [45], comprising approx. 2.5 million Java source files with 328 million lines of code. StoneDetector has been able to analyze this very large dataset in 79 hours. Again note that we report on an even faster version of StoneDetector when the Hamming distance is used as a string metric, as described in Sect. 4.7.
StoneDetector scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art.
4.6 Research Question 5: Configuration
SourcererCC NiCad CCAligner StoneDetector NIL Oreo
Recall 35,843 35,847 35,744 35,843 35,783 35,480
T 1 99.9% 99.9% 99.6% 99.9% 99.8% 98.9%
Recall 4,432 4,550 4,530 4,579 4,421 4,506
T 2 96.7% 99.3% 98.8% 99.9% 96.4% 98.3%
Recall 3,814 4,098 4,055 4,140 3,892 3,893
VST 3 91.6% 99.2% 97.4% 99.4% 93.5% 93.5%
Recall 9,026 10,680 10,534 13,008 10,083 12,625
ST 3 59.2% 70.0% 69.1% 85.3% 66.1% 82.8%
Recall 3,935 476 8,099 14,537 8,548 22,965
MT 3 4.8% 0.6% 10.0% 17.9% 10.5% 28.2%
Recall 1,690 11 12,673 11,612 14,500 52,839
T 4 0.0% 0.0% 0.2% 0.1% 0.2% 0.7%
Table 7: Recall of clone detectors on Java source code on BigCloneBench with a minimal code clone size of 6 lines of code (T 1 - Type 1, T 2 - Type 2, VST 3 - Very-Strongly Type 3, ST 3 - Strongly Type 3, MT 3 - Moderately Type 3, T 4 - Type 4)
Clone detection using StoneDetector should result in the highest possible recall with acceptable precision and runtime. In order to achieve this objective, certain parameters of our clone detection method have to be chosen. Thus, our fifth research question addresses experiments on optimal configuration parameters in StoneDetector, in particular considering the threshold values for syntactical similarity of code clones and their minimum size.
In the experiments presented so far, we have chosen 15 lines as the minimum size of code clone fragments. In order to justify our previous claim that smaller code fragments in general result in a skewed comparison, we performed additional experiments using different minimum sizes of code clones on BigCloneBench. In Table 7, we report recall values when a minimum size of 6 lines of code is used instead of 15 lines, since 6 lines of code are often used in the literature as well. Note that we have again omitted the tools Deckard and iClones due to their lower precision. As expected, the shown recall values are in general higher compared to those in Table 4. This can be well explained by Java methods with only 6 lines of source code, where half of the lines are just taken by the method’s annotations, signature, and curly braces. We, though, do not observe substantial changes in the tools’ comparisons, besides Oreo now outperforming StoneDetector in the case of Moderately Type 3 code clones. Furthermore, we also considered other minimum sizes in the range from 6 to 20 lines of code, though only for our tool StoneDetector. Due to space constraints, instead of providing the respective numbers, we have plotted the resulting precision and recall on BigCloneBench in Fig. 6. As can be seen, comparable recall values have been obtained, while precision values range from 96% to 100%, with increasing values for increasing minimum sizes of code clones. We can thus confirm our stated conclusion that increased minimum clone size supports the precision of clone detectors [10].
Figure 6: Additional experiments on parameter settings, considering StoneDetector’s used distance value $\tau$ (left) and admitted minimal size of code clones (right)
A second parameter we are interested in is the similarity threshold, which is used in our method to define whether two processed code fragments are considered code clones or not (cf. also Sect. 2.5). In the experiments conducted so far, this threshold value has been set to 70%, i.e., distance value $\tau=0.3$, which on the one-hand side is comparable to the threshold values as used in other clone detectors and on the other-hand side, in our experience, provides an optimal tradeoff of high recall and precision. In order to confirm this optimal tradeoff, we conduct additional experiments using StoneDetector on BigCloneBench using differing threshold values in terms of the distance value $\tau$ ranging from $\tau=0.1$ to $\tau=0.4$. The respective distance values are again plotted in conjunction with the resulting recall and precision in Fig. 6 (other configuration parameters equal those in Table 2, minimal size of code clones 15 lines of code). As expected, if the distance value increases, the recall increases with a simultaneous decrease in precision. The precision values range between 96% and 100%. It is striking that with low distance values, the recall values for Type 3 and Type 4 code clones are also lower. In these cases, the distance value is lower than the admitted syntactical difference of code clones. However, with increasing distance values, more code clones of Type 3 and Type 4 can be found again, though with a higher inaccuracy. As these results show, we can observe that recall and precision are balanced with a distance value of $\tau=0.3$, i.e., a similarity threshold of $70\%$.
On a more technical side, we conduct further experiments for providing parameter settings of StoneDetector for speeding up clone detection. First, in some cases, two code fragments can be ruled out as code clones by simply analyzing the structure of their respective description sets. For example, the factor in which the description sets of two code fragments differ may be too large, or the lengths of two respective paths in the code fragments’ description sets may deviate too much. In such cases, the comparison of the code fragments is short-circuited in StoneDetector for speeding up clone detection. In order to determine adequate parameter values, we run multiple experiments on BigCloneBench using differing values for the admitted maximal size factor of description sets and the admitted maximal length difference of paths in description sets. Based on these experiments, we can derive that a factor of 1.7 as the maximal size factor of description sets and a maximum length difference of up to 7 nodes in paths to be compared yields the best trade-off.
The abstraction of paths in StoneDetector, e.g., the uniform treatment of variables and constants as discussed in Sect. 2, can lead to several identical paths in one description set. As an example, consider two paths that only differ from each other in their last node, and where the respective nodes each represent assignments with which the same/different variables are assigned to different/the same constants. By abstracting variables and constants, these assignments are then mapped to the same representation, which results in an identical abstraction for both paths. We observed such presence of identical paths in one and the same description set to occur quite frequently. Although such occurrences do not influence StoneDetector’s recall or precision in finding code clones, its runtime is deteriorated due to the redundant comparisons of identical paths. Therefore, in additional experiments, we compare a variant of StoneDetector that does not handle identical paths and a variant where two identical paths or two paths that differ only in one node are merged into a single path. Surprisingly, this rather simple strategy already leads to a relative runtime improvement of 20%, which corresponds to a reduction from approx. 25 to 20 minutes on the whole BigCloneBench.
Clone detection on source code with StoneDetector achieves the best results using a similarity threshold of 70% ($\tau=0.3$). Increasing the minimal size of clones in general increases the precision of clone detectors.
4.7 Research Question 6: Metrics and Hashing
LCS with 4- LCSmodified Leven- Needleman- Hamming Hamming LCSmodified
byte hashing with MD5 shtein Wunsch w/o penalty with LSH
Recall 20,831 20,828 20,828 20,828 20,828 20,828 20,828
T 1 99.9% 99.9% 99.9% 99.9% 99.9% 99.9% 99.9%
Recall 3,476 3,475 3,475 3,475 3,475 3,475 3,475
T 2 100.0% 99.9% 99.9% 99.9% 99.9% 99.9% 99.9%
Recall 3,272 3,270 3,270 3,270 2,849 2,889 3,209
VST 3 99.4% 99.4% 99.4% 99.4% 86.6% 87.8% 97.6%
Recall 6,702 6,773 6,404 6,721 4,406 4,666 6,808
ST 3 90.5% 91.5% 86.4% 90.7% 59.5% 63.0% 92.0%
Recall 4,335 5,727 3,312 4,956 1,255 2,253 6,324
MT 3 16.1% 21.4% 12.3% 18.5% 4.4% 8.4% 23.6%
Recall 2,691 4,793 2,212 3,765 1,006 6,623 6,061
T 4 0.0% 0.1% 0.0% 0.0% 0.0% 0.1% 0.1%
#Clones 1,006,823 1,218,595 954,395 1,045,240 804,366 1,082,253 1,300,458
Precision 99% 99% 99% 99% 99% 97% 99%
Table 8: Recall of StoneDetector on Java source code on BigCloneBench using different string metrics and hashing methods (T 1 - Type 1, T 2 - Type 2, VST 3 - Very-Strongly Type 3, ST 3 - Strongly Type 3, MT 3 - Moderately Type 3, T 4 - Type 4)
Another parameter that can impact StoneDetector’s capability in finding code clones is the similarity measure, i.e., string metric, used for comparing the paths of two code fragments’ dominator trees (cf. also Sect. 2.3). Furthermore, since the path comparison is working on hashed fingerprints of the paths, the utilized hashing mechanism may as well influence StoneDetector’s performance. For our sixth research question, we therefore want to specifically analyze the role of the string metric and hashing mechanism. In our previously described experiments, we used a combination of a variant of the Longest Common Subsequence (LCS)-based metric ($\delta_{LCS_{modified}}$) in combination with Locality-Sensitive Hashing (LSH). This is in contrast to the authors’ previous work in [21], where we used the standard LCS-based metric in combination with a 4-byte prime number-based hashing mechanism.
In a first experiment, we considered the role of using different string metrics on clone detection in StoneDetector. To this end, we have chosen a subset of the string metrics implemented in StoneDetector, i.e., the LCS-based metric, modified LCS-based metric, Levenshtein distance, Needleman-Wunsch, Hamming distance, and modified Hamming distance. The latter metric depicts a variant of the Hamming distance in which different path lengths are not penalized, in the sense that the difference between different paths is not included in the distance measure as a penalty. We refer the interested reader to Sect. 2.3 for a short explanation of these string metrics. We analyze the role of a certain string metric by evaluating StoneDetector’s recall and precision on BigCloneBench when using the respective metric and MD5 hashing, keeping the other parameters stable (cf. Table 2). The results of this first experiment are shown in Table 8. Note that we did not observe any differences in the results when using either MD5 or a simple 4-byte hashing and are therefore denoting the results for the standard LCS-based metric with 4-byte hashing for the sake of comparison with our previous work [21]. In summary, the shown results indicate that using the Needleman-Wunsch algorithm or the modified LCS-based metric achieves the best performance for clone detection. Compared with these metrics, the use of the Levenshtein distance yields a visible deterioration in the recall values for Moderately and Strongly Type 3 code clones, while precision remains the same. In contrast, the Hamming distance implies significantly poorer recall values when looking at the syntactically more deviating code clones, i.e., clones of Type 3 and Type 4. While the use of the modified Hamming distance results in improved recall values, it features the worst precision. However, we observed significantly shorter runtimes of StoneDetector in the case of using either variant of the Hamming distance. Using the Hamming distance resulted in a speed-up of approximately 3.5 compared to the other metrics. As a rule of thumb to be taken from these measurements, we can thus state that the Hamming distance should be preferred over the other metrics when searching exact and near-miss clones, i.e., Type 1 and Type 2 code clones, due to its impact on shortening StoneDetector’s runtime.
In a second experiment, the hashing algorithm’s influence on StoneDetector’s performance is evaluated. Since the modified LCS-based metric has so far achieved the best results, we consider only this string metric in this experiment and compare the tool’s recall and precision when using either the MD5 hashing algorithm or Locality-Sensitive Hashing (LSH). Note again that we did not observe differences when using MD5 or a simple 4-byte hashing. As can also be seen in Table 8, the usage of LSH results in particular in improved recall values for Moderately Type 3 and Type 4 code clones, while maintaining precision. We have therefore chosen the modified LCS-based metric and LSH as StoneDetector’s default configuration in our evaluation.
Clone detection on source code with StoneDetector achieves the best results using the modified Longest Common Subsequence-based string metric in combination with Locality-Sensitive Hashing. Using the Hamming distance instead, though, allows for faster clone detection.
4.8 Research Question 7: Java Bytecode
StoneDetector (stack-based) StoneDetector (register-based) iClones (stack-based)
Recall 20,792 20,792 20,777
T 1 99.9% 99.9% 99.9%
Recall 3,463 3,466 3,373
T 2 99.7% 99.8% 97.2%
Recall 3,157 3,238 3,043
VST 3 97.2% 99.7% 93.7%
Recall 5,035 6,430 3,559
ST 3 68.9% 88.0% 48.7%
Recall 2,804 4,311 1,345
MT 3 10.9% 16.8% 5.2%
Recall 3,178 3,754 1,087
T 4 0.0% 0.0% 0.0%
#Clones 439,767 600,522 362,403
Runtime 77.20min 35.36min 37.59min
Precision 99% 99% 100%
Table 9: Recall of clone detectors on stack-/register-based Java Bytecode representations on BigCloneBench (T 1 - Type 1, T 2 - Type 2, VST 3 - Very-Strongly Type 3, ST 3 - Strongly Type 3, MT 3 - Moderately Type 3, T 4 - Type 4)
For our last research question, we consider Java’s Bytecode representation for clone detection. In general, we are interested in whether conventional code clone detectors like StoneDetector can be used for finding source code clones in their Java Bytecode representation. This use case may become apparent in situations where the source code is not available, e.g., when analyzing compiled Java code with regard to included malware.
Clone detection for Java Bytecode can be done either directly using the raw Bytecode or using an intermediate representation as generated, e.g., during compilation. The Soot framework [41] supports the transformation of Java Bytecode into several intermediate representations. Soot’s Baf format roughly resembles the original Bytecode and provides a stack-based intermediate representation. In contrast, the Jimple format is a register-based three-address code, which is generated by Soot out of the stack-based Bytecode for program analysis and transformation. We consider both formats in our evaluation and are interested in the their different suitability for code clone detection. Evaluating clone detectors for Java Bytecode is unfortunately not as straightforward as for Java source code, because the state-of-the-art benchmark BigCloneBench consists of only plain source files and is missing their respective dependencies [40]. We therefore use the Stubber tooling to transform BigCloneBench’s Java source files in such a way, that they can be compiled to Java Bytecode without the missing dependencies. For the details, we refer the interested reader to the publication introducing Stubber [47]. Using Stubber allows us to generate Bytecode for 52,816 or 95% of the overall 54,999 benchmark’s source files and to subsequently generate the Bytecode’s Baf and Jimple representation using Soot.
Due to characteristics of Java Bytecode as well as the Baf and Jimple formats, StoneDetector applies certain heuristics for eliminating some kind of low-level instructions that are specific to the instruction format. For example, with respect to the Baf format, there are several load and store instructions for pushing and pulling values onto and from the stack. As these instructions have no matching counterparts in the source code, they are eliminated before clone detection is performed in StoneDetector. Similar heuristics are applied to eliminate the duplicate paths in a dominator tree, which have been introduced in Jimple and Baf for the representation of the source code’s high-level control flow structures [40]. Furthermore, the LCS-based metric with 4-byte hashing is used, while code clones are again limited to code fragments with equal to or more than 15 lines of code. For the Baf format, $\tau=0.15$ is chosen, and for the Jimple format, again $\tau=0.3$ is used.
We also compare StoneDetector’s performance on Java Bytecode with another state-of-the-art clone detector, i.e., iClones. The tool does not readily come with support for Bytecode and we have thus extended iClones for the stack-based Java Bytecode. The results of our experiments on Java Bytecode are shown in Table 9. As can be seen, StoneDetector can be effectively used as a clone detector on Java Bytecode and provides a better performance with respect to recall and precision when compared to iClones. Note, though, that the latter tool has not been designed for Java Bytecode or other compiled code artifacts. We also note that StoneDetector performs better on the register-based Jimple format compared to the stack-based Baf format. This is justified by the fact that in the Bytecode, smaller changes in the original Java source code lead to larger differences in the intermediate format due to the characteristics of the instruction set and underlying stack machine model. While this also influences StoneDetector’s performance with Jimple, the deteriorating effects are larger for the Baf format. This plays no role with respect to exact or near-miss code clones of Type 1 and 2, as can also be seen in Table 9, but becomes more of an issue for Type 3 and 4 clones with a larger proportion of deviating source code instructions.
Detecting source code clones based on Java Bytecode is in general possible. A register-based representation of Java Bytecode is better suited for their detection compared to a stack-based representation.
5 Related Work
Techniques and tools for clone detection address the problem of finding code fragments that share a certain degree of syntactic and/or semantic similarity. Approaches for the detection of code clones can be classified according to the used detection methodology [6]. Following a recent survey on this research topic by Nasrabadi et al. [6], we differentiate text-, token-, tree-, graph-, and metric-based methods as well as hybrid, Deep Learning, and dynamic approaches. For an extensive discussion of code clone techniques and tools, we refer the interested reader to one of the various surveys available on the topic [6, 7, 1, 8] and here provide a condensed overview of the most formative and relevant work.
5.1 Text-based Techniques
The raw source code is used as input for the first group of methods. The source code is either used directly as a sequence of strings for finding matching parts of two code fragments, or only very limited pre-processing is applied beforehand, e.g., removal of whitespace and/or code normalization. Despite their simplicity and limited effectiveness on Type 3/4 code clones, text-based methods are still used today due to their efficiency and easy use for different programming languages. Examples include the general-purpose clone detectors like Simian [^17] https://simian.quandarypeak.com/ or plagiarism checkers like MOSS [48] [^18] https://theory.stanford.edu/~aiken/moss/.
As one of the first approaches, Baker proposed the tool Dup [49]. The tool combines a consistent identifier/literal renaming scheme with suffix trees on hashed code lines for finding parametrized code matches, i.e., near-miss code clones. Similarly, Johnson proposed to use the Karp-Rabin algorithm and sliding windows for hashed code lines, called fingerprints, to find duplicated code fragments [50]. In order to also detect Type 2 clones, he removes whitespace and comments and maps all identifiers, keywords, and literals to a common placeholder. The text-based clone detector Vuddy uses whole-function hashes on pre-processed source code for generating function fingerprints. These fingerprints are then used as indexes for the implementation of efficient lookups for vulnerable code in large codebases [51]. A well-known and up-to-date clone detector is NiCad [9, 52], which is mostly text-based but also applies lightweight parsing for normalizing the raw source code so that noise is removed and code formats are standardized. The code pre-processing can be adjusted to different programming languages using a customizable grammar definition. NiCad is using the Longest Common Subsequence algorithm for finding matching code and is one of the clone detectors used in our comparative evaluation in Sect. 4. The tool identifies almost all Type 1/2 clones and a substantial part of Type 3 code clones [12]. Various other authors have used NiCad as a starting point for their own research, e.g., Ragkhitwetsagul and Krinke [53] applied NiCad to compiled and subsequently decompiled code for further code normalization, while Chen et al. [54] used NiCad on decompiled Bytecode to identify Android malware.
5.2 Token-based Techniques
Token-based clone detectors use a lexer and/or parser for transforming raw source code into token sequences, which are then further analyzed to find matching code fragments. In this way, minor code edits, e.g., renaming of variables, can be easily handled. While token-based tools have in general shown good performance in the detection of near-miss code clones, more recent tools also provide good capabilities in the detection of large-gap Type 3 code clones featuring many code edits [10, 55, 56]. Furthermore, modern token-based clone detectors typically scale up to large code bases with several hundred million lines of code [18].
CCFinder [57] lexes two source code fragments into token sequences, thereby transforming some tokens according to certain rules, e.g., replacing mapping identifiers to a common token. The thus normalized token sequences are then handed over to the suffix tree algorithm for finding matching code fragments. Note that, as some of the above-mentioned text-based tools work quite similarly, the differentiation between text- and token-based approaches to clone detection is rather not strict. In subsequent work, the authors provided a flexible tokenization mechanism for easily extending the tool towards additional programming languages [58]. The suffix tree in its generalized form is also used by Göde and Koschke in their tool iClones for incremental code clone detection, i.e., detection of code clones in evolving code bases [17]. CP-Miner uses frequent subsequence mining for finding interleaved token subsequences in case of reordered statements or inserted code within code lines [36]. The token-based clone detector SourcererCC uses optimized indexing and filtering techniques for reducing the number of candidate code fragments when searching for code clones. In this way, the tool supports scalable clone detection for large code repositories with hundreds of millions of lines of code [18]. Scaling clone detectors for large codebases has also been the focus of subsequent research [59]. Another token-based technique is described by Ragkhitwetsagul and Krinke [60], their tool Siamese uses token sequences with different levels of abstractions and methods from document retrieval on their N-gram vector representations. As a result, they are able to provide a scalable method for the detection of Type 1, 2, and 3 code clones.
While tools like NiCad and SourcererCC are quite good at finding exact and near-miss clones, Type 3 and 4 clones with lower syntactical similarity provide a challenge. More recent token-based clone detectors, like CCAligner [10], LVMapper [56], and NIL [20], therefore focus on code clones with many consecutive code edits or many code modifications scattered around the code. These so-called large-gap clone detectors often apply more sophisticated string metrics for aligning and matching token sequences, e.g., e-edit distance [10] or Smith-Waterman sequence alignment [55]. We include the tools SourcererCC, iClones, and CCAligner as representative token-based clone detectors in our evaluation in Sect. 4
Approach
based on
Supported
clone types
Accuracy
Efficiency
Strength
Weakness
Text 1, 2, part of 3 + + language-agnostic fails on Type 3/4
Token 1, 2, part of 3 + + good scalability fails on Type 3/4
Tree/ 1, 2, 3, 4 + - uses syntactical and construction and
Graph semantic information matching expensive
Metric 1, 2, 3, 4 + $\circ$ language-agnostic threshold sensitivity
Deep 1, 2, 3, 4 + - fully automated, requires training,
learning very flexible prone to overfitting
Dynamic 1, 2, 3, 4 $\circ$ - finds semantic clones specialized for use case
Hybrid depends on technique
Table 10: Comparison of code clone detection techniques (- low, $\circ$ adequate, + high)
5.3 Tree- and graph-based Techniques
Tree- or graph-based techniques parse the source code into abstract syntax trees or use even further transformations for generating control flow or program dependence graphs. While abstract syntax trees contain all of the syntactical information about code fragments, edges in control flow graphs depict control dependences and in program dependence graphs control as well as data dependences and thus convey the semantics of code fragments. Thus, both approaches are able to detect code clones of Type 3 and Type 4, which share the same semantics but use differing syntax. The similarity matching for code fragments in tree- or graph-based techniques is performed on the generated program representations by finding subtrees or looking for graph isomorphisms. Both, the generation of the program representation as well as the tree or graph matching is often computationally expensive.
As one of the early tree-based clone detectors, CloneDR computes hash values for abstract syntax trees, which allows lowering the number of expensive tree comparisons, as only abstract syntax trees with the same hash value need to be compared [61]. More permissive tree comparisons are used by Evans et al. [62], where they apply, besides lexical abstractions, matching, e.g., differing identifiers, also structural abstractions, where whole subexpressions are matched, for finding more code clones on abstract syntax trees. Deckard is another tree-based tool, which uses the tree edit distance for matching code fragments based upon their representation as abstract syntax trees. In comparative evaluations, Deckard has been shown to find more Type 3/4 code clones, but at considerably lower precision [10, 11], a result we can reproduce in our evaluation in Sect. 4. Sager et al. use a combination of common subtree isomorphism and tree edit distance for detecting similar code fragments based upon a representation derived from abstract syntax trees [63]. Yang et al. linearize abstract syntax trees and apply the Smith-Waterman sequence alignment algorithm to the linearized trees for finding similar code fragments and thus code clones [64].
Komondoor and Horwitz present their clone detector PDG-DUP [65], where they use backward slicing on program dependence graphs for finding isomorphic subgraphs and thus code clones. While the slicing approach helps in increasing efficiency, the technique is still time-consuming for programs with large program dependence graphs. Krinke introduces duplix as another tool working with program dependence graphs, which iteratively searches for maximal similar subgraphs [66], but also notes the computational complexity of the approach. A similar approach based on isomorphic subgraph matching has been implemented for plagiarism detection [67]. Another approach, where slices on program dependence graphs are hashed and compared via MD5 or Locality-Sensitive Hashing (LSH) is implemented in the tool srcClone [32]. Breaking code into smaller code fragments, which are mapped to program dependence graphs, and using filtering based on different features of program dependence graphs before subgraph isomorphism testing allows CCSharp to lower the computation burden [68]. A similar multi-stage approach is described by Sargsyan et al. [69] using the LLVM framework for program dependence graph generation. Approximative graph matching using graph kernels is applied by the clone detector CCGraph [70] for speeding clone detection up based upon program dependence graphs. Other graph-based program representations besides program dependence graphs are rather used sporadically for clone detection, e.g., call graphs [71]. Using dominator trees for clone detection has been presented by the authors of this paper previously [72, 21]. The authors are not aware of any other research using dominator trees or dominator information for code clone detection.
5.4 Metric-based Techniques
A straightforward approach for finding similar code clones can be implemented by collecting characteristics of code fragments and matching code fragments with similar characteristics according to some variance threshold. Several metrics can be used for characterizing code fragments, including, e.g., the number of lines of code, the number of used variables and arguments, cyclomatic complexity, the number of various types of statements, and so on. While metric-based clone detectors are rather straightforward to implement and adjustable to different programming languages, they are quite sensitive to choosing the right threshold for similarity measurement. Further note that metric-based approaches can be seen as very simple data-driven approaches, since adding a mapping of the metrics into a vector space and an optimizable algorithm for distinguishing code clones from non-code clones based on this vector space yields the latter techniques [73]. In 1996, Mayrand et al. described a first approach for detecting code clones based upon a set of several metrics, including aspects like identifiers, layout, expressions, and control flow [74]. Similar function-level metrics, such as number of statements, McCabe’s cyclomatic complexity, and number of use-definition pairs, are used for clone detection by Patenaude et al. [75]. Metrics like type and order of statements, type and number of operators, and number of variables are also used by Sudhamani and Rangarajan [76] to build so-called feature tables for code fragments, which can then be compared using distance metrics like the city block metric. The technique showed reasonable results for Type 1, 2, and part of Type 3 clones. With the advent of multi-stage code clone detectors, metric-based techniques have also been proposed for pre-filtering candidate pairs of code fragments according to their metrics [11, 77].
5.5 Deep Learning Techniques
In the last decade, the application of machine learning and Deep Learning has become more widespread in clone detection. Common to these approaches is the embedding of code fragments into a manually crafted or automatically learned and often high-dimensional vector space, wherein statistical relationships are used for predicting similar code. The ability to thus automatically learn and recognize the hidden patterns of code clones from, e.g., today’s large and omnipresent code repositories, should guarantee promising results. While in particular Deep Learning clone detectors have shown very good performance in the detection of Type 1 to Type 4 code clones lately, research has recently also highlighted common problems and pitfalls in the application of Deep Learning to clone detection, i.e., training/evaluation bias, overfitting, and missing ground truth [78, 79, 80].
Prior approaches considered conventional machine learning methods. For instance, Cesare et al. use algorithms like naive Bayes and random forests for detecting package-level clones [81]. Clustering via k-Means can also be used for unsupervised learning of similar code and for finding code clones with respect to various similarity thresholds [82]. More recent work, however, has shifted towards the application of deep neural networks. In pioneering research, White et al. use an autoencoder architecture for unsupervised learning of abstract syntax trees and compute code similarities on the learned representation [83]. The Deep Learning clone detector CCLearner uses the frequencies of token categories in the abstract syntax trees of code fragments as features and learns to discriminate code clones from non-code clones using these features via supervised training of a deep neural network [84]. CDLH [85] is using the long short-term memory (LSTM) architecture for representation learning on abstract syntax trees and is trained to embed the resulting representation to hash values, such that code clones are close in terms of their Hamming distance. Similarly, Zhang et al. are using, in their architecture ASTNN, bidirectional recurrent units on embedded statement trees for encoding abstract syntax trees and identifying code clones [86]. Different settings for embeddings of abstract syntax trees are evaluated by Büch and Andrzejak [87], where they use a Siamese architecture that optimizes the embeddings of two abstract syntax trees such that the cosine similarity in the case of code clones is maximized. As an interesting result, pre-trained embeddings like Word2vec, Code2vec [24] have proven useful. Another lightweight approach for embedding abstract syntax trees is presented by Gao et al. [88], where the authors use random walks on the abstract syntax trees and embed the resulting token sequences similar to Word2vec. The resulting encodings of code fragments can then be compared via their Euclidean distances.
The Siamese architecture is also used in the hybrid clone detector Oreo, where the model is used with metrics-based feature vectors for detecting harder-to-find clones, achieving in particular good results for Moderately Type 3 clones and beyond [11]. We use Oreo as a representative for Deep Learning clone detectors in our evaluation in Sect. 4.
The more specialized Deep Learning architecture FA-AST-GMN for the detection of code clones based on graph matching networks [89] has been introduced for abstract syntax trees [90]. Furthermore, other code representations besides abstract syntax trees have also been considered as input for Deep Learning clone detectors [23, 91]. Control and data flow is combined into a rich feature model in DeepSim for training a Deep Learning model for clone detection [91]. Fang et al. combine abstract syntax trees and control flow graphs in so-called fusion embeddings for finding in particular Type 4 code clones [92]. Patel and Sinah [93] combine this approach with the Siamese architecture. The usage of the different code representations: identifiers, abstract syntax trees, control flow graphs, and Bytecode, as well as their combined usage for clone detection is further studied by Tufano et al. [94]. Various code metrics and embeddings of API calls have also been used as input features to a Siamese network for cross-language clone detection [73]. Finally, pre-trained models based on the Transformer architecture, e.g., CodeBERT [95], GraphCodeBERT [96], and even large language models like GPT [97] have been evaluated and adjusted for clone detection [98, 99]. Note that this paper focuses on conventional approaches to code clone detection, which the authors believe to play a role in the future. We, though, acknowledge the potential of such methods and are interested in the in-depth analysis of their strengths and weaknesses in future work.
5.6 Dynamic Techniques
Dynamic methods focus on the computations of code instead of its syntax and usually target binary code. Therefore, dynamic analyses or testing is used to execute code fragments with sample inputs and to track their behavior. The resulting data can then be used to match semantically similar code and, in this way, in particular, to detect Type 4 code clones. As for all dynamic methods, scalability and execution time can be problematic, and accuracy is tightly coupled to the coverage of test inputs.
Pewny et al. identify semantic code clones in binary code by first translating the binary code into a common intermediate representation, which is then interpreted block-wise for random input values. The resulting input/output pairs of code fragments are hashed and the semantic hashes can afterward be used for finding semantically similar code [100]. Similar approaches are chosen by Hu et al. [101], where behavior of code fragments is emulated, and by Jiang and Su [102], where code fragments are instrumented and executed on sample inputs for tracking input/output behavior. While the above-mentioned approaches consider compiled C/C++ binaries, other techniques have been proposed for Java as well [103, 104]. Instrumentation of binary code is also used by Su et al. [105] for generating dynamic dependency graphs, where semantically code fragments are discovered via searching isomorphic subgraphs. Alternatively, matching semantic clones can be done done using the notion of abstract memory states, which are derived by a static analyzer and compared based on the abstract states’ semantics [106]. In this way, clone detection is again aligned with the semantics of code fragments and independent of syntactical similarity. Due to the different focus, i.e., heterologously instead of homologously developed code, and specialized use case of these dynamic approaches, we do not include them in our evaluation.
5.7 Hybrid Techniques
Hybrid techniques combine the aforementioned approaches for aggregating their advantages. For example, as can also be seen in Table 10, textual or token-based techniques usually suffer from lower effectivity in identifying Type 3 and Type 4 code clones with larger syntactical variance, and tree- or graph-based techniques demonstrate increased effectivity in finding harder-to-detect code clones but are mostly less efficient. Thus, a combination of both should provide an effective and efficient code clone detector. Besides, we frequently also found combinations of text- and token-based techniques as well as token- or tree-based techniques with Deep Learning approaches, i.e., data-driven techniques. Since distinguishing the conventional approaches and hybrid approaches cannot, as a matter of fact, be strict, some hybrid techniques have already been discussed before. For example, Nicad [9, 52] combines a basically textual approach to clone detection with lightweight and token-based preprocessing, Yang et al. generate the abstract syntax tree of code fragments for encoding structural information but then compare the linearized trees via the token-based sequence alignment [64], and Oreo [11] uses a two-step approach by using a metric-based technique as a basis for finding code clones and a Deep Learning model for the remaining harder-to-detect code clones. In the following, we present some other representative examples which fit into the category of hybrid approaches.
Token-based and tree- or graph-based approaches to clone detection are often used in such a way that a token-based technique is used for filtering code clone candidates which are afterward validated using the tree-/graph-based technique [107]. A combination of token- and tree-based clone detection is also used by Vislavski et al. [108] in their tool LICCA to detect cross-language code clones and by Wang et al. [109] in their tool CCStokener to speed up detection efficiency while retaining detection performance. Similar two-step pipelines have been presented for metric-based approaches, where the metric-based technique is used for prefiltering clone candidates and a graph- or text-based technique is used for validation [110, 111].
Further research focused on the role of different intermediate representations for code clone detection, e.g., considering Java source code and Bytecode. Selim et al. apply the clone detector CCFinder [57] on the Jimple representation of Java Bytecode and find source code clones that are not detected otherwise. They thus propose to use their approach to complement clone detectors working solely on the source code [112]. A similar method has been described for the usage of NiCad [9] on LLVM-based intermediate representations for C [113]. A more comprehensive approach, including a study of token-, tree-, and graph- based code representations and a combination thereof within the tool TACC is presented by Wang et al. [114], where they demonstrate an increased performance for the combined approach for all types of code clones. The usage of different code representations in combination with Deep Learning is also analyzed by Tufano et al. [94].
6 Conclusion
In this paper, we have presented an approach to code clone detection for the Java programming language using dominator trees and approximate string matching. Using dominator trees allows us to take advantage of control flow information in order to find harder-to-detect code clones with many syntactical deviations or semantic clones, which feature low syntactical similarity but share the same computations. The clone matching itself is, though, based upon string metrics, which are applied to encoded paths of the dominator trees for exploiting the advantages of conventional clone detecting, in particular in the case of exact and near-miss code clones. In this way, our method is different from most other conventional approaches to clone detection for Java, since the approach incorporates control flow besides syntax and, as a result, achieves better performance when asked to find code clones with larger syntactical variance, i.e., Type 3 and Type 4 clones.
The method for clone detection has been implemented in the publicly available StoneDetector platform, which provides a versatile system for finding code clones in both Java source code and Bytecode, as well as supporting manifold experiments on clone detection’s parameters. Using the implementation in StoneDetector, we conducted a thorough evaluation of our approach with respect to the state-of-the-art benchmarks for Java code clone detection, i.e., BigCloneBench, Google Code Jam, GPTCloneBench, and Project CodeNet. Furthermore, we experimentally compared the performance of the code clone detection approach in StoneDetector to other state-of-the-art clone detectors, including NiCad, iClones, CCAligner, SourcererCC, Deckard, Oreo, and NIL. The experiments proved StoneDetector’s competitive capabilities in finding exact near-miss clones with respect to precision and recall. In addition, StoneDetector excels at identifying code clones with larger syntactical variance at high precision when compared to the other state-of-the-art tools. Further experiments demonstrated StoneDetector’s ability to analyze large code bases with up to multiple million lines of code. Besides its performance, we highlight the versatility of the platform for experimenting with different parameters for code clone detection, including the similarity threshold, minimal size of code clones, string metrics, e.g., Hamming distance, Longest Common Subsequence, and hashing methods, e.g., Locality-Sensitive Hashing. Eventually, StoneDetector’s ready-made extensibility to other programming languages and program representations is presented, as is shown for stack-based and register-based Bytecode formats.
While this paper focuses on conventional code clone detection and therefore omits an in-depth discussion and comparative evaluation with Deep Learning approaches to code clone detection, we want to analyze and incorporate such models into the StoneDetector platform in our future work. While we acknowledge the promising results of, in particular, large language models in recent work [97], as discussed in the related work section, we also assume that conventional methods of clone detection conjoined with their advantages, as presented in this paper, will play a role in the future. We though concede that the assessment of the advantages and disadvantages of conventional code clone detection in comparison with Deep Learning approaches, as well as their incorporation into a possible combined approach, requires further research and analysis. Besides, we are also interested in extending the StoneDetector platform with respect to other programming languages, which have not been in the research focus for code clone detection.

### v2

@@
 Since such code clones, i.e., identical or similar fragments of code, can bloat software projects and cause issues like bug or vulnerability propagation, their identification is of importance.
-In this paper, we present the StoneDetector platform and its underlying method for finding code clones in Java source and Bytecode.
+In this paper, we present StoneDetector and its underlying method for finding code clones in Java source and Bytecode.
 StoneDetector implements a conventional clone detection approach based upon the textual comparison of paths derived from the code’s representation by dominator trees.
 In this way, the tool does not only find exact and syntactically similar near-miss code clones, but also code clones that are harder to detect due to their larger variety in the syntax.
-We demonstrate StoneDetector’s versatility as a conventional clone detection platform and analyze its various available configuration parameters, including the usage of different string metrics, hashing algorithms, etc.
+We demonstrate StoneDetector’s versatility as a conventional clone detection tool and analyze its various available configuration parameters, including the usage of different string metrics, hashing algorithms, etc.
 In our exhaustive evaluation with other conventional clone detectors on several state-of-the-art benchmarks, we can show StoneDetector’s performance and scalability in finding code clones in both, Java source and Bytecode.
+keywords:
+code clone , clone detection , Java , static analysis
+journal: Journal of Systems and Software
 […]
+[dhge]organization=Cooperative University Gera-Eisenach,addressline=Weg der Freundschaft 4, city=Gera, postcode=07546, country=Germany
 […]
+[fsu]organization=Friedrich Schiller University Jena,addressline=Ernst-Abbe-Platz 2, city=Jena, postcode=07743, country=Germany
 1 Introduction
-[…]
 int func(int[] array) {
@@
 Code clones [83] are identical or similar fragments of code that, e.g., arise due to copy-and-paste practices in software development.
-Since such practices are quite common, code clones can be frequently found, as empirical studies on code duplication have shown in the past [2, 3].
+Since such practices are quite common, code clones can be frequently found, as empirical studies on code duplication have shown in the past [84, 123, 64].
 While code cloning may initially speed up software development or allow for reusing working code, it also introduces long-term challenges: Duplicated code inflates code bases and increases the maintenance burden, as developers have to manually track and update all copies for fixing bugs or implementing code changes.
@@
 In this paper, we address this problem and are interested in advancing the detection of code clones with larger syntactical variance.
-[…]
 1private static void zipDir(File zipDir, ZipOutputStream zos, String path) {
@@
 26}
-[…]
 1public static void zipDir(File zipDir, ZipOutputStream zos, String path) {
@@
 As a consequence, the method is able to provide better performance in the detection of code clones with larger syntactical variance, i.e., Type 3 and Type 4 code clones.
-The presented clone detection method is implemented in the StoneDetector platform, which is publicly available [^1] https://stonedetector.fmi.uni-jena.de/ and allows for finding code clones in Java source and Bytecode.
-Using the platform enabled us to perform an in-depth evaluation of the method’s performance on various state-of-the-art benchmarks, including BigCloneBench[12, 13], Google Code Jam, Project CodeNet [15], and GPTCloneBench [16].
-As a result, we found that our clone detection method and its implementation in the StoneDetector platform allow for finding exact and near-miss code clones in Java source code at competitive performance with other conventional code clone detectors, including NiCad [14], iClones [17], SourcererCC [18], Deckard [19], CCAligner [10], NIL [20], and Oreo [11].
+The presented clone detection method is implemented in StoneDetector, which is publicly available [^1] https://stonedetector.fmi.uni-jena.de/ and allows for finding code clones in Java source and Bytecode.
+Using StoneDetector enabled us to perform an in-depth evaluation of the method’s performance on various state-of-the-art benchmarks, including BigCloneBench [102, 103], Google Code Jam, Project CodeNet [78], and GPTCloneBench [4].
+As a result, we found that our clone detection method and its implementation in StoneDetector allow for finding exact and near-miss code clones in Java source code at competitive performance with other conventional code clone detectors, including NiCad [22], iClones [31], SourcererCC [88], Deckard [39], CCAligner [110], NIL [69], CloneWorks [104] and Oreo [87].
 Additionally, StoneDetector in particular excels at finding source code clones with larger syntactical variance at high precision.
 StoneDetector also scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art.
-Furthermore, we have experimented with different configuration parameters for clone detection, e.g., threshold value, code clone size, applied string metrics, and hashing algorithms, etc., and thereby show the versatility of the StoneDetector platform besides identifying its optimal configuration.
+Furthermore, we have experimented with different configuration parameters for clone detection, e.g., threshold value, code clone size, applied string metrics, and hashing algorithms, etc., and thereby show the versatility of StoneDetector besides identifying its optimal configuration.
 Eventually, we analyzed the method’s potential in also finding code clones in compilation artifacts like Java Bytecode, in cases where the source code is unavailable in the first place.
@@
   • 1.
-    We demonstrate the feasibility of finding code clones in Java code fragments by encoding and textually comparing paths of the code fragments’ dominator trees.
+    We demonstrate the feasibility of finding code clones in Java source code fragments by encoding and textually comparing paths of the code fragments’ dominator trees.
   • 2.
-    We provide the publicly available StoneDetector platform with an implementation of our code clone detection method as a versatile conventional code clone detector and evaluate its performance and scalability.
+    We provide the publicly available tool StoneDetector with an implementation of our code clone detection method as a versatile conventional code clone detector and evaluate its performance and scalability.
   • 3.
@@
 Furthermore, expanding the approach for detecting large-gap and subclones, as presented in our previous conference paper [8], we added locality-sensitive hashing (LSH) and a modified LCS metric to StoneDetector, which resulted in performance gains for these types of code clones.
-We now also include a more exhaustive analysis of StoneDetector’s performance in finding harder-to-detect Type 3/4 code clones by including three more state-of-the-art benchmarks for code clone detection besides BigCloneBench, i.e., Google Code Jam, Project CodeNet, and GPTCloneBench, and two more clone detectors, i.e., iClones and NIL, in the evaluation.
+We now also include a more exhaustive analysis of StoneDetector’s performance in finding harder-to-detect Type 3/4 code clones by including three more state-of-the-art benchmarks for code clone detection besides BigCloneBench, i.e., Google Code Jam, Project CodeNet, and GPTCloneBench, and three more clone detectors, i.e., iClones, CloneWorks [104], and NIL, in the evaluation.
 Moreover, we also extended our previous approach in Amme et al.
-[21] for clone detection on stack-based and register-based Java Bytecode representations, such that the StoneDetector platform can be used for both finding code clones in Java source code as well as Java Bytecode.
-The rest of the paper is structured as follows: Our clone detection approach and its underlying code clone detection method based on encodings of paths in the dominator tree and string metrics for path comparison are described in detail in Sect.
-2.
-Sect.
-3 introduces the StoneDetector platform’s architecture and implementation.
+[8] for clone detection on stack-based and register-based Java Bytecode representations, such that StoneDetector can be used for both finding code clones in Java source code and Bytecode.
+The rest of the paper is structured as follows: Sect.
+2 introduces StoneDetector’s overall approach and architecture.
+The underlying clone detection method based on encodings of paths in the dominator tree and string metrics for path comparison are described in detail in Sect.
+3.
 In Sect.
-4, we present the in-depth evaluation of our approach on various benchmarks for Java code clone detection, including the comparison with state-of-the-art clone detectors and the thorough analysis of the StoneDetector platform’s various configuration parameters.
+4, we present the in-depth evaluation of our approach on various benchmarks for Java code clone detection, including the comparison with state-of-the-art clone detectors and the thorough analysis of StoneDetector’s various configuration parameters.
 Related work on code clone detection is discussed in Sect.
 5 and Sect.
-6 finally concludes the paper.
+6 finally concludes the paper and mentions future work.
+2 Overall Approach and Architecture
+Our approach to code clone detection has been implemented in StoneDetector [8, 93] thus providing a full-scale stand-alone clone detector for Java.
+As input, StoneDetector expects a number of Java source or Bytecode files for which the tool generates a list of all code clone pairs, which are identified and mapped back to the original code by file location and line numbers.
+The resulting information can then be used, for example, to search for similar programs in code repositories [47], to analyze the evolution of code [72], or to track bugs and vulnerabilities propagated via duplicated code [61].
+Figure 3: Overview of the tool StoneDetector
+An overview of StoneDetector can be seen in Fig.
+3, which shows its inner structure and how its components interact with each other.
+Considering, to begin with, Java source code, the tool starts with performing lexical and syntax analysis on the given Java source files.
+This step yields the abstract syntax trees, which are subsequently subject to control flow analysis for generating (abstract) dominator trees.
+Afterward, paths are extracted from the generated dominator trees and linearized into description sets.
+The individual nodes of the description sets can be further encoded using a hashing algorithm, which yields fingerprint sets.
+Either raw description sets or fingerprint sets are eventually transferred to the clone matcher, which checks for similarities using one out of the several available string metrics.
+Based upon this similarity analysis, code clones are identified and returned for further processing.
+We will next provide more information on the individual components of the described processing pipeline for Java source code files.
+For a more elaborative introduction of the clone detection method based on dominator tree and string metrics we refer the reader to Sect.
+3.
+Lexical/Syntax Analysis
+The lexical and syntax analysis in the Java source code frontend of StoneDetector is utilizing the Spoon analysis framework [76] [^2] https://spoon.gforge.inria.fr/.
+The Spoon framework includes libraries for parsing Java source code files up to Java version 20 into abstract syntax trees (AST) and provides the AST metamodel and functionality for manipulating them.
+Control/Dominator Analysis
+The Spoon framework also supports generating control flow graphs of individual Java functions based on its AST model.
+We can therefore resort to Spoon’s control flow builder for our control flow analysis.
+However, the framework currently includes neither a meta-model for dominator trees nor any dominator analysis.
+For the construction of dominator trees, the respective builder from the WALA framework [24] [^3] https://github.com/wala/WALA has been integrated into StoneDetector.
+As before, dominator trees are built at the function level.
+In addition, due to the coarse abstraction for Java’s exception handling in control flow graphs generated by Spoon, we have implemented a standard interprocedural exception analysis for Java as described by Chang et al.
+[17].
+Using the analysis allows for a more fine-grained control flow mapping of Java’s exception handlin and experiments have shown that enabling exception analysis implies better results for clone detection.
+Path Extraction and Description Set Encoding
+Dominator trees are transformed into description sets in order to subsequently perform textual clone matching.
+To this end, a (abstract) dominator tree is traversed such that the set of its paths starting at the root node and ending at a leaf nodes, respectively, is generated.
+Note that we in general perform identifier and literal abstraction in the resulting description sets, much like as in other clone detectors, e.g., NiCad [85].
+Though function names at call sites may optionally be conserved, which provided better experimental results in some cases.
+We therefore offer a switch for enabling/disabling identifier abstraction for function names at call sites.
+Likewise, StoneDetector treats encoding of description sets into fingerprint sets via a hashing function as optional.
+Raw description sets can thus be directly committed to the clone matcher or beforehand transformed such that the contained instructions are encoded using one of the several implemented hashing algorithms, i.e., MD5, 4-byte prime number hashing, or Locality-Sensitive Hashing (LSH).
+Clone Matching
+The clone matching component receives a set of description or fingerprint sets as input and performs pairwise similarity checks.
+The similarity of description sets, i.e., code fragments, is traced back to the accumulated similarity of its paths in our approach.
+Path similarity is computed using one out of several string metrics available in StoneDetector.
+Among the currently available metrics are the Hamming distance [34], the Levenshtein distance [56], the Needleman-Wunsch-based metric [71], as well as the LCS-based metric [12].
+Further supported metrics include the N-Gram, Jaccard, Cosine, Jaco-Winkler, and Damerau-Levenshtein metrics.
+The thus defined processing pipeline in StoneDetector is parallelized.
+On the one hand, generation of description sets can be done in parallel with the help of streams.
+If all description sets have been eventually generated, clone matching begins.
+This process, on the other hand, can also be done in parallel, again using streams.
+The configuration of a specific run of StoneDetector allows for the specification of the number of threads to be used in both computations.
+In order to further increase throughput, especially when processing larger numbers of code fragments, a copy strategy can optionally be used for description set handling.
+NiCad uses a similar technique for improving its runtime behavior [85, 22].
+In StoneDetector, however, the technique has been adopted in such a way that the number of identified clones remains unchanged whether or not the technique is enabled.
+The basic idea of the copy strategy is to continue with further comparisons only for one of the two code fragments $f$ and $g$ whenever a clone pair $f$ and $g$ is found: Suppose comparisons are continued for fragment $f$ and terminated for fragment $g$.
+After all comparisons are finished, code clones found for $f$ are added to the clones already found for $g$, up to termination.
+Note that in our variant of the copy strategy, two code fragments are only considered as a clone pair if they are identical to each other, i.e., if they have a distance value of zero.
+Using our copy strategy, runtime behavior does not improve as much as in NiCad, however, accuracy of the resulting code clones is guaranteed.
+Apparently, as can be seen by the above discussion, StoneDetector provides a rich set of configuration parameters that can be used for optimizing clone detection and demonstrates the tool’s versatility.
+Among the most critical parameters are the threshold value $\tau$ defining the required similarity of code clones, the minimum size $l$ of code fragments acceptable as code clones, the string metric $m$ used for path matching, and the hash function h used for description set encoding (cf.
+also Table 1).
+The therein given default values of configuration parameters are chosen for the tool’s best tradeoff of recall and precision with acceptable runtime.
+For an in-depth description of the StoneDetector’s configuration parameters, we refer the reader to the evaluation in Sect.
+4 and to the supplementary information available online [^4] https://stonedetector.fmi.uni-jena.de/.
+Configuration parameter Value range Default value
+Similarity threshold $\tau$ $0\leq\tau\leq 1.0$ $0.3$
+Minimum clone size in code lines $l$ $0\leq l$ 15
+String metric $\delta$ LCS LCSmodified
+LCSmodified
+Levenshtein
+Needleman-Wunsch
+Hamming
+N-Gram
+Jaccard
+Cosinus
+Jaco-Winkler
+Damerau-Levenshtein
+Hashing algorithm $h$ none LSH
+4-byte hashing
+MD5
+LSH
+Table 1: Main configuration parameters in StoneDetector
+2.1 Extensibility of StoneDetector
+The use of fingerprints to represent description sets makes clone detection in StoneDetector fundamentally independent of the selected programming language and possible program representation.
+In fact, as is also depicted in Fig.
+3, StoneDetector can be divided into a frontend and backend, with the former generating description sets of given code fragments and the latter performing the actual clone matching.
+The strict separation of frontend and backend allows us to facilitate a readily extensible tool, which can be easily extended for processing other programming languages as well as other internal program representations, such as control flow graphs, abstract syntax trees, or even just token sequences of linear code.
+Such an extension is only required to map the new programming language or program representations to fingerprints by using a certain hash function.
+Clone matching can subsequently be performed using the existing backend.
+The specific hash function used for construction is irrelevant to the functionality of the backend.
+In this way, fingerprints form a general and elegant interface for extending our system to other programming languages and program representations.
+To demonstrate the StoneDetector’s versatility and extensibility, we have added the handling of Java Bytecode as another frontend of StoneDetector [94] (cf.
+Fig.
+3).
+The new frontend gets a set of class files as input and transforms them to fingerprint sets, which are committed to the existing clone matcher of StoneDetector.
+Specifically, processing class files is implemented via the Soot framework [107], which first transforms the contained Bytecode into the Bytecode-like Baf or into the register-based Jimple representation.
+Afterward, dominator trees are generated based upon the Baf or Jimple representation and paths contained in the dominator trees are extracted and encoded into fingerprint sets, as required for clone matching.
 3 Clone Detection using Dominator Trees
@@
 2, i.e., function zipDir, can be found in Fig.
-LABEL:fig2 (a).
+4 (a).
+(a) Control Flow Graph
+(b) Dominator Tree
+(c) Abstract Dominator Tree
+Dset= [[=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
+       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[CALL, #3, V, V, L],
+       [=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
+       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[=, V, NEW,  #4, V]->[=, V, NEW, #5]->
+       [CALL, #6, V]->[COND, NE, =, V, CALL, #7, V, NEG, V]->[CALL, #8, V, V, V],
+       [=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
+       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[=, V, NEW,  #4, V]->[=, V, NEW, #5]->
+       [CALL, #6, V]->[COND, NE, =, V, CALL, #7, V, NEG, V]->[CALL, #9],
+       [=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
+       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[POSTINC, V]]
+Fingerprints=
+[35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
+ c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
+ a45688ae3fc84165d19f2efe859e43db->c088b1671e63a009662f8e31c773a0a0,
+ 35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
+ c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
+ a45688ae3fc84165d19f2efe859e43db->c21875dc52243cb4f19c52d3e02be40a->19b4a2a27bfa0ccbeeee5564571ec011->
+ cb4b446eaed074420c4e9c663ffd3741->c15ffb85b338cca72f6729adcc304469,
+ 35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
+ c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
+ a45688ae3fc84165d19f2efe859e43db->055bfda33fe11d9fb52471e283c27d9c->19b4a2a27bfa0ccbeeee5564571ec011->
+ cb4b446eaed074420c4e9c663ffd3741->f13c29192ec234b0a51eb4f32bc85efe,
+ 35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
+ c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
+ a45688ae3fc84165d19f2efe859e43db->fc6ebb514a9a13b92cff0ae97894c23e]
+(d) Description Set and Fingerprints
+Figure 4: Control flow graph, dominator tree, description set, and fingerprints of example
 A dominator tree [1] is a tree and models a certain control flow property of code fragments, specifically, which instruction precedes the other under all possible executions of the respective code fragment.
@@
 An example of a dominator tree is visualized in Fig.
-LABEL:fig2 (b).
+4 (b).
 As can be seen, the shown dominator tree corresponds to the control flow graph in Fig.
-LABEL:fig2(a).
+4(a).
 Note that in the dominator tree, each node is reachable from the start node, i.e., the tree’s root, via a single unambiguous path.
@@
 We use an abstract, text-oriented encoding of instructions for our clone detection method, yielding the abstract dominator tree which is depicted in Fig.
-LABEL:fig2 (c).
+4 (c).
 Therein, instructions and their respective operands are denoted in preorder notation.
@@
 Fig.
-LABEL:fig2).
+4).
 Function calls are optionally represented using a constant pool.
-A constant pool access is denoted by the character # and the respective function’s entry in the constant pool, as, e.g., in [CALL,#6,V] for instruction zos.putNextEntry(anEntry) (again, cf.
+A constant pool access is denoted by the character # and the respective function’s entry in the constant pool, as, e.g., in in [CALL,#6,V] for instruction zos.putNextEntry(anEntry) (again, cf.
 Fig.
-LABEL:fig2).
+4).
 By using abstract dominator trees, we are able to more easily match code fragments that do not completely align with each other syntactically, in particular in the case of Type 2 and part of Type 3/4 code clone deviations.
@@
 [6], where they use a linearization of abstract syntax trees based on paths that start and end at distinct leaf nodes of an abstract syntax tree.
-Though, in contrast, in our method, given a dominator tree $D=(N,E^{*})$, we consider all paths starting at the tree’s root node $s\in N$ and ending at a respective leaf node $l\in\{n\in N\,|\,\not\exists(n,m)\in E^{*}\}$.
+Though, in contrast, in our method, given a dominator tree $D=(N,E^{*})$, we consider all paths starting at the tree’s root node $s\in N$ and ending at a respective leaf node $l\in\{n\in N\,|\,\nexists(n,m)\in E^{*}\}$.
 The resulting set of these paths can then be used to unambiguously represent the dominator tree $D$.
@@
 For example, considering the abstract dominator tree in Fig.
-LABEL:fig2 (c), the path from the start node to the rightmost leaf node is encoded by the sequence:
+4 (c), the path from the start node to the rightmost leaf node is encoded by the sequence:
 […]
@@
 In the case of the abstract dominator tree depicted in Fig.
-LABEL:fig2 (c), we have four paths starting at the start node and ending in the leftmost, the left inner, the right inner, and the rightmost leaf node, respectively.
+4 (c), we have four paths starting at the start node and ending in the leftmost, the left inner, the right inner, and the rightmost leaf node, respectively.
 Accordingly, the description set shown in Fig.
-LABEL:fig2 (d) contains the four instruction sequences representing the paths.
+4 (d) contains the four instruction sequences representing the paths.
 Given two dominator trees $D$ and $D^{\prime}$, which include two code fragments, $f$ and $g$, respectively, we call $f$ and $g$ strictly similar iff both description sets $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$ have the same number of elements and there exists a similar path $p^{\prime}\in D^{\prime}_{\mathit{set}}$ for each path $p\in D_{\mathit{set}}$.
@@
 Note that the conventional Hamming distance therefore tends to penalize paths with different lengths.
+Note that the Hamming distance can be computed quite efficiently, which makes it a good candidate for fast and scalable near-miss code clone detection, as we have also observed in our experiments (cf.
+Sect.
+4).
 Example:
   • Code fragment 1: for (int j=0; j<<m; j++) { acc=acc-j; proc(acc); }
-  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
+  • $f=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[=,V,+,V,L]}$
   • Code fragment 2: for (int l=0; l<<m; l++) { acc=acc+l; }
-  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow$ [POSTINC,V]
+  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow$ [=,V,+,V,L]
 Hamming:
@@
     Substitute [=,V,+,V,V] with [=,V,-,V,V]
-    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[POSTINC,V]}$
+    $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$
   • 2.
-    Substitute [POSTINC,V] with [CALL,#4,V]
-    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$
-  • Hamming distance $\delta_{\mathit{H}}(p,q)=3=2+1$ because $p$ has one more character than $q$.
+    Substitute [=,V,+,V,L] with [CALL,#4,V]
+    $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$
+  • Hamming distance $\delta_{\mathit{H}}(f,g)=3=2+1$ because $f$ has one more character than $g$.
 Levenshtein:
@@
     Substitute [=,V,+,V,V] with [=,V,-,V,V]
-    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[POSTINC,V]}$
+    $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$
   • 2.
     Insert [CALL,#4,V]
-  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[\mbox{CALL},\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
-  • Levenshtein distance $\delta_{\mathit{L}}(p,q)=2$.
+  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[\mbox{CALL},\#4,V]}\rightarrow\texttt{[=,V,+,V,L]}$
+  • Levenshtein distance $\delta_{\mathit{L}}(f,g)=2$.
 Needleman-Wunsch:
-  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$ $\rightarrow\texttt{[POSTINC,V]}$
-  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[POSTINC,V]}$
-  • Needleman-Wunsch-based distance $\delta_{\mathit{NW}}(p,q)=max(|p|-score(p,q),|q|-score(p,q))=2$, where $score(p,q)=3$.
+  • $f=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$ $\rightarrow\texttt{[=,V,+,V,L]}$
+  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,+,V,L]}$
+  • Needleman-Wunsch-based distance $\delta_{\mathit{NW}}(f,g)=max(|f|-score(f,g),|g|-score(f,g))=2$, where $score(f,g)=3$.
 Longest Common Subsequence (LCS):
-  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
-  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\ \rightarrow\texttt{[POSTINC,V]}$
-  • $z=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[POSTINC,V]}$
-  • LCS-based distance $\delta_{\mathit{LCS}}(f,g)=max(|p|-|z|,|q|-|z|)=2$.
+  • $f=\texttt{[=,V,L]}\rightarrow\texttt{[\textless,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[=,V,+,V,L]}$
+  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[\textless,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\ \rightarrow\texttt{[=,V,+,V,L]}$
+  • $z=\texttt{[=,V,L]}\rightarrow\texttt{[\textless,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$
+  • LCS-based distance $\delta_{\mathit{LCS}}(f,g)=max(|f|-|z|,|g|-|z|)=2$.
 Figure 5: Distance metrics for calculating path similarity
@@
 In the second step, the optimal string alignment with maximal total score is computed by performing a backward traversal on this similarity matrix.
+While the Levenshtein and Needleman-Wunsch-based distance require more computational effort, they are also more permissive in case of code clones with larger syntactical variance as in case of large-gap clones, where matching code fragments have many and/or larger edits similar to DNA sequence alignments [63].
 The application of string metrics to the comparison of paths in abstract dominator trees is illustrated in Fig.
 5 by way of an example.
-In the upper part, two code fragments are shown, conjoined with two paths $p$ and $q$ from their abstract dominator trees.
+In the upper part, two code fragments are shown, conjoined with two paths $f$ and $g$ from their abstract dominator trees.
 As can be seen below, the (abstract) instructions take on the role of characters in the visualized calculations of string metrics.
 In the alignment provided for the Needleman-Wunsch algorithm, vertical lines between nodes represent matches and horizontal lines represent inserted gaps.
-For the value of $store(p,q)$, a scheme is used in which the weights for matches are set to $1$, for mismatches to 1-1, and for gaps to 0.
+For the value of $score(f,g)$, a scheme is used in which the weights for matches are set to $1$, for mismatches to 1-1, and for gaps to 0.
 The Longest Common Subsequence (LCS) algorithm [12, 85] is a well-known tool for comparing the contents of files, used for example in the Unix diff utility or in the revision control system git.
 Besides, clone detectors like NiCad are relying on the LCS algorithm as well.
-In our case, the algorithm is applied to paths instead of lines of code, such that the longest common subpath is computed, e.g., $\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[POSTINC,V]}$ for paths $p$ and $q$ in Fig.
-4.
+In our case, the algorithm is applied to paths instead of lines of code, such that the longest common subpath is computed, e.g., $\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$ for paths $f$ and $g$ in Fig.
+5.
 This longest common subpath then forms the basis for defining the metric $\delta$ for two paths of dominator trees, which equals the maximum number of unique nodes for both paths.
 Using the original LCS-based metric, though, can become problematic in cases of large-gap code clones [110], e.g., code clones where one code fragment is embedded into the other.
-The LCS metric then has issues in recognizing the similarity of the respective paths of the code fragments’ dominator trees, since the paths differ too much in their length.
-For this reason, a modified variant of the LCS-based metric can be used, which performs additional similarity checking in situations where two paths are not considered similar by the original LCS-based metric and analyzes whether or not the derived LCS distance value is close to the length of a respective smaller path.
-On a technical note, code fragments consisting of only one basic block and thus path are then considered as code clones if the calculated LCS value corresponds to a value lower than 30% of the number of nodes of the smaller code fragment.
-Furthermore, a path $p$ is also considered a subpath of a path $q$ if the calculated LCS value of $p$ and $q$ is less than 10% of the length of $p$.
-Finally, the smaller of both paths is required to have a minimum length of 10 nodes.
-Note that there exist other string metrics as well, besides the ones discussed above, which have also been implemented in the StoneDetector platform.
-Among these are, e.g., N-gram distance, cosine distance, Jaccard similarity.
+The LCS metric then has problems recognizing the similarity of the respective paths of the dominator trees of the code fragments, as the paths differ too greatly in length.
+For this reason, a modified variant of the LCS-based metric can be used, which performs an additional similarity check in situations where two paths are not considered similar by the original LCS-based metric and therefore analyzes whether the length of the derived longest common subsequence is close to the length of the respective smaller path.
+If this is the case, the unique nodes in the smaller path are considered for building the LCS metric measure, rather than those of the larger path.
+Technically speaking, code fragments that consist of only one basic block are considered code clones if the calculated LCS value corresponds to a value that is less than a percentage of the number of nodes of the smaller code fragment that is specified by a certain threshold.
+This additional detection feature for subclones can also be applied to such code fragments that do not consist of only one basic block.
+For that purpose, for two subpaths $q$ and $p$, for which the length of the longest common subsequence is close to the number of nodes of the smaller path $p$, the LCS value is determined using the uniquely occurring nodes in $p$.
+However, to avoid overly imprecise clone analysis, the examination for subpaths only makes sense if the length of the longest common subsequence of the two paths is nearby to the length of the smaller path $p$ and, furthermore, a minimum length is assumed for $p$.
+In our later described evaluation experiments (cf.
+Sect.
+4.5), we observed that a maximum length difference of 10% in the number of nodes in the longest common subsequence and $p$ and a minimum length of 10 lines for $p$ is a good practical choice for these parameters.
+Note that there exist other string metrics as well, besides the ones discussed above, which have also been implemented in StoneDetector.
+Among these are, e.g., N-gram distance, cosine distance, Jaccard similarity, etc.
 For the sake of brevity, we here limit ourselves to the subset of metrics that we use in the evaluation in Sect.
@@
 For example, the MD5 algorithm [82] provides a widely used hash function with 16-byte hash values, which is often used as a checksum of file downloads and suffices for creating collision-free encodings of instructions in our abstract dominator trees.
-The MD5 algorithm is the default hashing algorithm for fingerprints in the StoneDetector platform, and its application for encoding the description sets of the example in Fig.
-LABEL:fig2 is shown in the lower part of the same figure.
+The MD5 algorithm is the default hashing algorithm for fingerprints in StoneDetector, and its application for encoding the description sets of the example in Fig.
+4 is shown in the lower part of the same figure.
 We have implemented and experimented with two other hashing algorithms as well, including a simple 4-byte hash function based upon prime number encoding and Locality-Sensitive Hashing.
@@
 3.3 and Sect.
-2.4, for the distance function $\delta$ allows us to compare description sets and subsequently assess the similarity of code fragments.
+3.4, for the distance function $\delta$ allows us to compare description sets and subsequently assess the similarity of code fragments.
 To this end, given two code fragments with respective description sets $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$, the most similar path $q$ of $D^{\prime}_{\mathit{set}}$ is identified for each path $p$ of $D_{\mathit{set}}$, such that $q$ has a minimal value $\delta(p,q)$, and the arithmetic mean for these values over all paths $p\in D_{\mathit{set}}$ is computed.
@@
 Using $\tau$, we have $\Delta(D_{\mathit{set}},D^{\prime}_{\mathit{set}})<\tau$ for two description sets $D_{\mathit{set}},D^{\prime}_{\mathit{set}}\in\mathbb{D_{\mathit{set}}}$ iff $D_{\mathit{set}},D^{\prime}_{\mathit{set}}$ are similar and their corresponding code fragments constitute a code clone.
-3 Architecture and Implementation
-The approach to code clone detection introduced in Sect.
-2 has been implemented in the StoneDetector platform [21, 33] thus providing a full-scale stand-alone clone detector for Java.
-As input, StoneDetector expects a number of Java source or Bytecode files for which the platform generates a list of all code clone pairs, which are identified and mapped back to the original code by file location and line numbers.
-The resulting information can then be used, for example, to search for similar programs in code repositories [34], to analyze the evolution of code [35], or to track bugs and vulnerabilities propagated via duplicated code [36].
-Figure 5: Overview of the StoneDetector platform
-An overview of the StoneDetector platform can be seen in Fig.
-5, which shows its inner structure and how its components interact with each other.
-Considering, to begin with, Java source code, the platform starts with performing lexical and syntax analysis on the given Java source files.
-This step yields the abstract syntax trees, which are subsequently subject to control flow analysis for generating (abstract) dominator trees.
-Afterward, paths are extracted from the generated dominator trees and linearized into description sets.
-In case fingerprints are enabled, the individual nodes of the description sets are further encoded using a hashing algorithm, which yields fingerprint sets.
-Either raw description sets or fingerprint sets are eventually transferred to the clone matcher, which checks for similarities using one out of the several available string metrics.
-Based upon this similarity analysis, code clones are identified and returned for further processing.
-We will next provide further details on the individual components of the described processing pipeline for Java source code files:
-Lexical/Syntax Analysis
-The lexical and syntax analysis in the Java source code frontend of the StoneDetector platform is utilizing the Spoon analysis framework [37] [^2] https://spoon.gforge.inria.fr/.
-The Spoon framework includes libraries for parsing Java source code files up to Java version 20 into abstract syntax trees (AST) and provides the AST metamodel and functionality for manipulating them.
-Control Flow/Dominator Analysis
-The Spoon framework also supports generating control flow graphs of individual Java functions based on its AST model.
-We can therefore resort to Spoon’s control flow builder for our control flow analysis.
-However, the framework currently includes neither a meta-model for dominator trees nor any dominator analysis.
-For the construction of dominator trees, the respective builder from the WALA framework [38] [^3] https://github.com/wala/WALA has been integrated into the StoneDetector platform.
-As before, dominator trees are built at the function level.
-In addition, due to the coarse abstraction for Java’s exception handling in control flow graphs generated by Spoon, we have implemented a standard interprocedural exception analysis for Java as described by Chang et al.
-[39].
-Using the analysis allows for a more fine-grained control flow mapping of Java’s exception handling.
-The analysis is, though, optional, but experiments have shown that enabling exception analysis implies better results for clone detection.
-Path Extraction and Description Set Encoding
-As described in Sect.
-2, dominator trees are transformed into description sets in order to subsequently perform textual clone matching.
-To this end, a (abstract) dominator tree is traversed such that the set of its paths starting at the root node and ending at leaf nodes, respectively, is generated (cf.
-Fig.
-LABEL:fig2).
-Note that we in general perform identifier and literal abstraction in the resulting description sets, much like as in other clone detectors, e.g., NiCad [9].
-Though function names at call sites may optionally be conserved, which provided better experimental results in some cases.
-We therefore offer a switch for enabling/disabling identifier abstraction for function names at call sites.
-Likewise, the StoneDetector platform treats encoding of description sets into fingerprint sets via a hashing function as optional.
-Raw description sets can thus be directly committed to the clone matcher or beforehand transformed such that the contained instructions are encoded using one of the several implemented hashing algorithms, i.e., MD5, 4-byte prime number hashing, or LSH.
-Clone Matching
-The clone matching component receives a set of description or fingerprint sets as input and performs pairwise similarity checks.
-As described in Sect.
-2, the similarity of description sets, i.e., code fragments, is traced back to the accumulated similarity of its paths in our approach.
-Path similarity is computed using one out of several string metrics available in the StoneDetector platform.
-Among the currently available metrics are the Hamming distance [25], both the normal and the weighted Levenshtein distance [26], the Needleman-Wunsch-based metric [27], as well as the original or modified LCS-based metric [28].
-Further supported metrics include the N-Gram, Jaccard, cosine, Jaco-Winkler, and Damerau-Levenshtein metrics.
-The thus defined processing pipeline in the StoneDetector platform is parallelized.
-On the one hand, generation of description sets can be done in parallel with the help of streams.
-If all description sets have been eventually generated, clone matching begins.
-This process, on the other hand, can also be done in parallel, again using streams.
-The configuration of a specific run of the StoneDetector platform allows for the specification of the number of threads to be used in both computations.
-In order to further increase throughput, especially when processing larger numbers of code fragments, a copy strategy can optionally be used for description set handling.
-NiCad uses a similar technique for improving its runtime behavior [9, 14].
-In StoneDetector, however, the technique has been adopted in such a way that the number of identified clones remains unchanged whether or not the technique is enabled.
-The basic idea of the copy strategy is to continue with further comparisons only for one of the two code fragments $f$ and $g$ whenever a clone pair $f$ and $g$ is found: Suppose comparisons are continued for fragment $f$ and terminated for fragment $g$.
-After all comparisons are finished, code clones found for $f$ are added to the clones already found for $g$, up to termination.
-Note that in contrast to NiCad’s implementation, in our variant of the copy strategy, two code fragments are only considered as a clone pair if they are identical to each other, i.e., if they have a distance value of zero.
-Using our copy strategy, runtime behavior does not improve as much as in NiCad, however, accuracy of the resulting code clones is guaranteed.
-Apparently, as can be seen by the above discussion, the StoneDetector platform provides a rich set of configuration parameters that can be used for optimizing clone detection and demonstrates the platform’s versatility.
-Among the most critical parameters are the threshold value $\tau$ defining the required similarity of code clones, the minimum size $l$ of code fragments acceptable as code clones, the string metric $m$ used for path matching, and the hash function h used for description set encoding (cf.
-also Table 1).
-For an in-depth description of the platform’s configuration parameters, we refer the reader to the supplementary information available online [^4] https://stonedetector.fmi.uni-jena.de/.
-Configuration parameter Value range Default value
-Similarity threshold $\tau$ $0\leq\tau\leq 1.0$ $0.3$
-Minimum clone size in code lines $l$ $0\leq l$ 15
-String metric $\delta$ LCS LCS
-LCSmodified
-Levenshtein
-Needleman-Wunsch
-Hamming
-N-Gram
-Jaccard
-Cosinus
-Jaco-Winkler
-Damerau-Levenshtein
-Hashing algorithm $h$ none MD5
-4-byte hashing
-MD5
-LSH
-Table 1: Main configuration parameters in the StoneDetector platform
-3.1 Extensibility of the StoneDetector platform
-The use of fingerprints to represent description sets makes clone detection in the StoneDetector platform fundamentally independent of the selected programming language and possible program representation.
-In fact, as is also depicted in Fig.
-5, StoneDetector can be divided into a frontend and backend, with the former generating description sets of given code fragments and the latter performing the actual clone matching.
-The strict separation of frontend and backend allows us to facilitate a readily extensible platform, which can be easily extended for processing other programming languages as well as other internal program representations, such as control flow graphs, abstract syntax trees, or even just token sequences of linear code.
-Such an extension is only required to map the new programming language or program representation to fingerprints by using a certain hash function.
-The clone matching can subsequently be performed using the existing backend of the StoneDetector platform.
-The specific hash function used for construction is irrelevant to the functionality of the backend.
-In this way, fingerprints form a general and elegant interface for extending our system to a wide variety of programming languages and program representations.
-To demonstrate the platform’s versatility and extensibility, we have added the handling of Java Bytecode as another frontend of the StoneDetector platform [40] (cf.
-Fig.
-5).
-The new frontend gets a set of class files as input and transforms them to fingerprint sets, which are committed to the existing clone matcher of StoneDetector.
-Specifically, processing class files is implemented via the Soot framework [41], which first transforms the contained Bytecode into the Bytecode-like Baf or into the register-based Jimple representation.
-Afterward, dominator trees are generated based upon the Baf or Jimple representation and paths contained in the dominator trees are extracted and encoded into fingerprint sets, as required for clone matching.
 4 Evaluation
 In the evaluation, we are interested in finding answers to a number of research questions, in particular about the performance of our code clone detector compared to the current state of the art.
-We also present results that we believe to provide general insights into code clone detection for Java and to help other researchers when implementing code clone detectors, besides showing the versatility of the StoneDetector platform.
+We also present results that we believe to provide general insights into code clone detection for Java and to help other researchers when implementing code clone detectors, besides showing the versatility of StoneDetector.
 Thus, we address the following research questions:
   • 1.
-    Research Question 1: How many code clones does our clone detection method and tool StoneDetector detect, i.e., what is its recall compared to the state of the art?
+    Research Question 1: How many code clones does our clone detection method and tool StoneDetector detect, i.e., what is its recall compared to the state of the art, and how many of StoneDetector’s reported code clones are real code clones, i.e., what is the tool’s precision compared to the state of the art?
   • 2.
-    Research Question 2: How many of StoneDetector’s reported code clones are real code clones, i.e., what is the tool’s precision compared to the state of the art?
+    Research Question 2: Is the conventional code clone detector StoneDetector in general able to find more difficult-to-identify code clones, i.e., code clones with larger syntactical deviation?
   • 3.
-    Research Question 3: Are conventional code clone detectors like StoneDetector in general able to find more difficult-to-identify code clones, i.e., code clones with larger syntactical deviation?
+    Research Question 3: Does StoneDetector scale to large codebases containing up to multiple million lines of code?
   • 4.
-    Research Question 4: Does StoneDetector scale to large codebases containing up to multiple million lines of code?
+    Research Question 4: How do the used threshold value and the minimal size of code clone influence StoneDetector’s performance? Which role do different string metrics and hashing methods play?
   • 5.
-    Research Question 5: How do the used threshold value and the minimal size of code clone influence StoneDetector’s performance?
-  • 6.
-    Research Question 6: Which role do different string metrics and hashing methods play in a code clone detector like StoneDetector?
-  • 7.
-    Research Question 7: Can source code clones also be found in Java Bytecode using clone detectors like StoneDetector?
+    Research Question 5: Can source code clones also be found in Java Bytecode using clone detectors like StoneDetector?
 Tool Configuration parameters
@@
 70% threshold
+CloneWorks 70% similarity threshold, aggressive configuration:
+line splitting, blind renaming, literal abstraction
 Table 2: Tools’ configuration parameters
@@
 However, in order to avoid an evaluation skewed to this benchmark and to assess the performance of clone detectors for the more difficult-to-identify code clones of Type 3 and Type 4, we also consider the benchmarks Google Code Jam [^6] https://zibada.guru/gcj/, Project CodeNet [78] [^7] https://github.com/IBM/Project_CodeNet, and GPTCloneBench [4] [^8] https://github.com/srlabUsask/GPTCloneBench.
-In the comparison of StoneDetector with other code clone detectors for Java, a set of representative state-of-the-art tools has been chosen, including conventional text-based, token-based, tree-based, and data-driven clone detectors, namely, we use the tools:
+In the comparison of StoneDetector with other code clone detectors for Java, a set of representative state-of-the-art tools has been chosen, including conventional text-based, token-based, tree and graph-based, as well as hybrid clone detectors (cf.
+Sect.
+5).
+While we focus on conventional code clone detectors in this paper, we have also added the Oreo hybrid tool, which also implements a Deep Learning approach, for the sake of comparison.
+Namely, we use the following tools:
 Benchmark Type 1 Type 2 Very-Strongly Strongly Moderatly Type 4
@@
 Sect.
-4.6).
+4.5).
 The clone detectors’ configurations are otherwise chosen in conformance with the literature; see also Table 2.
+In case of StoneDetector, the tool’ configuration has been chosen for the tool’s best tradeoff of recall and precision with acceptable runtime.
+When deviating from this default configuration, we will explain it in the respective sections.
 All experiments have been performed on a Linux workstation with an Intel Core i7-10700K processor and 32 GB RAM.
@@
 Note that the benchmark BigCloneBench is quite skewed towards harder-to-find Type 4 code clones, as can be seen in Table 3.
+The usage of BigCloneBench as dataset for harder-to-detect code clones of Type 3 and Type 4 have been controversially discussed in recent years[58, 53].
+In particular, its naive usage as dataset for Deep learning clone detectors can compromise training and evaluation [92, 52].
+The benchmark suffers from several issues, which are caused by its construction.
+For example, the benchmark is not only skewed towards Type 4 code clones but also towards certain functionalities, where, e.g., the Copy file functionality holds half of all code clones [53].
+Besides, its ground truth quality, i.e., true code clones labels, is disputed for Type 4 code clones [58, 53].
+However, note that these issues do not impair its usage for evaluating clone detectors on Type-1, Type-2, and Type-3 clones.
+However, in order to underpin our findings for harder-to-detect code clones of Type 3 and Type 4, we augment our evaluation with other state-of-the-art benchmarks for code clone detection.
 A benchmark including in particular harder-to-detect code clones has been defined with GPTCloneBench [4].
 The authors of the benchmark used the code clones of the SemanticBench [3] benchmark in prompts for generating additional code clones with the GPT-3 large language model.
-Manually validating the generated code clones afterward resulted in a set of 37,14 code clones, of which 16,989 are clone pairs with Java source code that we can use in our experiments.
+Manually validating the generated code clones afterward resulted in a set of 37,149 code clones, of which 16,989 are clone pairs with Java source code that we can use in our experiments.
 In spite of the benchmark’s authors intentions, we use the 5,431 near-miss code clones of Type 1 and Type 2 as well as the 11,558 Type 3 and Type 4 code clones in the experiments.
@@
 We consider solutions for the year 2022, and, as only part of the solutions are Java code, we are able thus to use 10,936,986 code clone pairs in our experiments.
+Note that Google Code Jam also suffers from being biased towards certain dominant functionalities, i.e., programming tasks [58].
 The last benchmark also originates from programming contests.
@@
 Around 5% of the code fragments, or 25 million lines of code, in Project CodeNet are Java code, which results in 11,212,500 code clone pairs used in our experiments.
-Note that all three aforementioned benchmarks, in contrast to BigCloneBench, do not further distinguish between the different clone types.
-Deckard SourcererCC NiCad iClones CCAligner StoneDetector NIL Oreo
-Recall 18,318 20,827 20,829 20,842 20,831 20,828 20,844 20,844
-T 1 87.8% 99.9% 99.9% 99.9% 99.9% 99.9% 100% 100%
-Recall 3,317 3,444 3,460 3,453 3,456 3,475 3,460 3,475
-T 2 95.4% 99.1% 99.5% 99.3% 99.4% 99.9% 99.5% 99.9%
-Recall 2,139 3,209 3,243 3,253 3,218 3,209 3,271 3,283
-VST 3 65.0% 97.6% 98.6% 98.9% 97.8% 97.6% 99.5% 99.8%
-Recall 1,806 5,784 6,246 6,531 5,943 6,808 6,549 6,584
-ST 3 24.4% 78.1% 84.3% 88.2% 80.2% 92.0% 88.5% 88.9%
-Recall 2,311 1,462 169 4,051 1,784 6,324 3,499 5,187
-MT 3 8.6% 5.5% 0.6% 15.1% 6.6% 23.6% 13.1% 19.4%
-Recall 57,063 250 7 6,554 2,133 6,061 3,821 6,821
-T 4 1.3% 0.0% 0.0% 0.1% 0.0% 0.1% 0.1% 0.1%
-#Clones 13,239,074 804,315 754,022 987,659 1,244,160 1,300,458 1,274,852 1,015,425
-Precision 46% 100% 100% 97% 98% 99% 99% 99%
+Apparently, in case of Google Code Jam and Project CodeNet, code clone samples from programming contests may not be representative of production code [58].
+We though believe that using a wide range of benchmarks allows for a generalizable interpretation of our evaluation results, even if we are aware of the benchmarks’ issues.
+Note that all three benchmarks, in contrast to BigCloneBench, do not distinguish between the different clone types.
+Deckard SourcererCC NiCad iClones CCAligner StoneDetector NIL CloneWorks Oreo
+Recall 18,318 20,827 20,829 20,842 20,831 20,828 20,844 20,828 20,844
+T 1 87.8% 99.9% 99.9% 99.9% 99.9% 99.9% 100% 99.9% 100%
+Recall 3,317 3,444 3,460 3,453 3,456 3,475 3,460 3,460 3,475
+T 2 95.4% 99.1% 99.5% 99.3% 99.4% 99.9% 99.5% 99.5% 99.9%
+Recall 2,139 3,209 3,243 3,253 3,218 3,209 3,271 3,240 3,283
+VST 3 65.0% 97.6% 98.6% 98.9% 97.8% 97.6% 99.5% 98.5% 99.8%
+Recall 1,806 5,784 6,246 6,531 5,943 6,808 6,549 7,073 6,584
+ST 3 24.4% 78.1% 84.3% 88.2% 80.2% 92.0% 88.5% 95.5% 88.9%
+Recall 2,311 1,462 169 4,051 1,784 6,324 3,499 1,305 5,187
+MT 3 8.6% 5.5% 0.6% 15.1% 6.6% 23.6% 13.1% 4.8% 19.4%
+Recall 57,063 250 7 6,554 2,133 6,061 3,821 19 6,821
+T 4 1.3% 0.0% 0.0% 0.1% 0.0% 0.1% 0.1% 0.0% 0.1%
+#Clones 13,239,074 804,315 754,022 987,659 1,244,160 1,300,458 1,274,852 915,434 1,015,425
+Precision 46% 100% 100% 97% 98% 99% 99% 100% 99%
 Table 4: Recall and precision of clone detectors on Java source code on BigCloneBench (T 1 - Type 1, T 2 - Type 2, VST 3 - Very-Strongly Type 3, ST 3 - Strongly Type 3, MT 3 - Moderately Type 3, T 4 - Type 4, #Clones - total number of detected clones)
-4.2 Research Question 1: Recall
+4.2 Research Question 1: Recall and Precision
 In order to address the first research question about how many code clones in Java source code our clone detection method is able to find, we ran StoneDetector on BigCloneBench via BigCloneEval [103].
 For a comparison with the state of the art, we also ran seven other clone detectors on the same benchmark.
-The resulting numbers of identified known code clones, i.e., recall values, for the tools Deckard, SourcererCC, NiCad, iClones, CCAligner, StoneDetector, NIL, and Oreo are shown in Table 4, as measured by BigCloneEval.
+The resulting number of identified known code clones, i.e., recall values, for the tools Deckard, SourcererCC, NiCad, iClones, CCAligner, StoneDetector, NIL, CloneWorks, and Oreo are shown in Table 4, as measured by BigCloneEval.
 Since these numbers reflect only the annotated code clones in the benchmark (cf.
-also Sect.
+Sect.
 4.1), we additionally list the overall number of detected clones.
 As noted before, we only consider clones with a minimum size of 15 lines of code.
-Duplicate clones, as reported by some of the tools, have been removed.
+Duplicate clones, as reported by some tools, were removed.
 Recall values are given in absolute and relative numbers.
@@
 Thus, we can state that StoneDetector performs competitively with the state of the art for exact and near-miss code clones.
+Note though that in case of Type 1 clones, i.e., exact duplicates of code fragments that only differ in layout, comments, or formatting (cf.
+Sect 1), StoneDetector misses certain code clones.
+Closer inspection revealed that these cases can be attributed to a low number of code samples from the benchmark which can not be properly processed or parsed using the Spoon analysis framework.
 As expected, beginning with code clones of Strongly Type 3, recall values are declining due to the larger syntactical differences of code fragments.
-We can also observe a difference between near-miss code clone detectors, e.g., NiCad and SourcererCC, and tools addressing harder-to-find clones, i.e., CCAligner, iClones, Oreo, StoneDetector, and NIL.
+We can also observe a difference between near-miss code clone detectors, e.g., NiCad, CloneWorks, and SourcererCC, and tools addressing harder-to-find clones, i.e., CCAligner, iClones, Oreo, StoneDetector, and NIL.
 StoneDetector achieves the best recall values for code clones of Strongly Type 3 and Moderately Type 3, but is overtaken by the tools Deckard, iClones, and Oreo for code clones of Type 4.
 However, the high number of Type 4 clones Deckard is finding has to be taken with care, considering its poor precision as described in the following section and also reported in the literature [88, 110].
-When compared with state-of-the-art conventional clone detectors for Java source code, StoneDetector detects code clones with competitive recall and finds more code clones with 30% to 50% syntactical variance.
-4.3 Research Question 2: Precision
 In addition to a high number of identified code clones, a code clone detector shall also avoid reporting on wrong code clones.
 Thus, we have to measure the clone detector’s precision, i.e., how many of the reported code clones are actual code clones.
-Unfortunately, answering the second research question is challenging, since BigCloneBench/BigCloneEval only contains information about a subset of all possible code clones in the benchmark.
+Unfortunately, answering this question is challenging, since BigCloneBench/BigCloneEval only contains information about a subset of all possible code clones in the benchmark.
 Due to the benchmark’s construction, only code clones of the considered 43 functionalities are known [102].
 We have therefore performed a manual evaluation on random samples of the tool’s reported code clones, as is also usually conducted in the respective research literature [110, 87, 88].
-To this end, we randomly selected samples of 400 clone pairs of the overall reported code clones for each clone detector.
+To this end, we randomly selected samples of 400 clone pairs of the overall code reported clones for each clone detector.
 The clone pairs were blindly, i.e., without revealing the respective tool, given to the three authors of this paper, who acted as judges.
@@
 [110].
-StoneDetector detects code clones in Java source code, similar to other state-of-the-art clone detectors, at very high precision.
+When compared with state-of-the-art conventional clone detectors for Java source code, StoneDetector detects code clones with competitive recall and finds more code clones with 30% to 50% syntactical variance.
+At the same time, StoneDetector detects code clones in Java source code, similar to other state-of-the-art clone detectors, at very high precision.
+4.2.1 Qualitative Analysis
 We conducted further studies on the characteristics of the code clones found by StoneDetector and the other clone detectors.
-To this end, we made pairwise comparisons of the detected code clones for StoneDetector and each other tool, respectively, and looked into their common and exclusively identified code clones.
-On the one hand, we were able to observe a large fraction of commonly identified code clones.
+To this end, we first made pairwise comparisons of the detected code clones for StoneDetector and some of the other tools, respectively, and looked into their common and exclusively identified code clones.
+On the one hand, we were thereby able to observe a large fraction of commonly identified code clones.
 This mostly points to the subset of exact and near-miss clones as identified by all clone detectors (cf.
 also Table 4).
-On the other hand, we also found a significant portion of code clones with larger syntactical variance, in particular structural code clones and subclones, that are exclusively identified by StoneDetector and not by the other tools.
-For more details, we refer the reader to our previous study in [21].
-4.4 Research Question 3: Generalizability
-Our third research question addresses the capability of clone detectors to identify code clones in source code with larger syntactical variability, in particular code clones of Strongly Type 3 and beyond.
-Besides, we are also interested in analyzing the clone detectors’ performance on benchmarks other than BigCloneBench in order to prevent the evaluation from being biased toward this particular benchmark.
-We therefore ran the clone detectors SourcererCC, NiCad, CCAligner, StoneDetector, NIL, and Oreo on the three benchmarks GPTCloneBench, Google Code Jam, and Project CodeNet, respectively, and report the observed recall values in Table 5.
+On the other hand, we also found a significant portion of code clones with larger syntactical variance, in particular structural code clones, i.e, code clones which implement the same control flow using different syntactical constructs, and subclones, i.e., code clones where the control flow of one code fragment is embedded into the other fragment’s control flow [8].
+Code clones of these types are often exclusively identified by StoneDetector and not by the other tools.
+When analyzing clone pairs that are detected by StoneDetector but not by the other tools, we created random samples of found code clones on BigCloneBench, specifically for comparing StoneDetector with NiCad, Oreo, CCAligner, and CloneWorks, as representatives of near-miss clone detectors, tools focusing on large-gap clones, and Deep Learning approaches.
+For this purpose, for each tool $A$, we generated a random sample of 100 clones for the set of code clones identified by StoneDetector but not by AA ($StoneDetector\setminus A$) and for the set of code clones detected by $A$ but not by StoneDetector ($A\setminus StoneDetector$), respectively, and manually inspected those samples.
+Finally, we also created a random sample of 100 code clones which were only detected by StoneDetector and not by any other tool considered in our evaluation, though excluding Deckard due to its low precision (i.e., NiCad, SourcererCC, CCAligner, iClones, NIL, Oreo, and CloneWorks).
+StoneDetector vs NiCad
+The closer examination of 100 random clone pairs that were detected by StoneDetector but not by NiCad revealed that StoneDetector finds more clones with greater syntactic variance and in particular, structural clones (55) and subclones (35).
+The remaining clones were classified as of other types (4) and false positives (6).
+Analysis of 100 clones found only by NiCad revealed that these were mainly due to NiCad’s pretty printing, which often leads to an increase in code size and thus higher permissiveness for code clones, and due to different path sizes in the dominator trees in StoneDetector and token sequences in NiCad.
+StoneDetector vs CloneWorks
+In CloneWorks, clone detection is based on a modified Jaccard metric.
+This metric represents code fragments as sets of contained tokens and calculates their minimum rate of overlapping tokens [104].
+Similar to Oreo, CloneWorks thus does not relies on the order of code fragment’s instructions for clone detection.
+As a result, CloneWorks can also detect simple structural clones and subclones, if the lengths of the code fragments do not differ too much.
+Our examination of clone pairs found exclusively by StoneDetector showed that more structural clones (28) and subclones (64) were detected.
+A closer inspection revealed that the clones often differed greatly in the number of lines of code.
+Further clones found exclusively by StoneDetector were classified as of other clone types (6) and false positives (2).
+Clone pairs found exclusively by CloneWorks can be attributed to the use of token sets.
+Structural clones (24) and subclones (49) were consequently also found by CloneWorks if the order of the instructions in the clone pairs differs.
+StoneDetector cannot detect such clones because its used LCS method requires a sequence of instructions.
+For example, code fragments with a high number of identical instructions occurring at different code locations cause issues for LCS and are often not recognized as clones even though they contain the very same instructions.
+Other code clones found exclusively by CloneWorks were such clones, in which instructions appeared in short form in one code fragment and in full form in the other (7), e.g., as in case of a foreach loop and a normal for loop over iterators.
+In contrast to StoneDetector, which conceives the latter loop using four sub-expressions, CloneWorks treats both loops as one expression in its aggressive version.
+Remaining code clones found exclusively by CloneWorks were again of others clone types (17) and false positives (3).
+StoneDetector vs CCAligner
+Our manual inspection of code clones found by StoneDetector but not by CCAligner revealed that CCAligner misses many subclones, which are detected by StoneDetector (52), which can be attributed to its code window approach failing in case of scattered code modifications.
+Furthermore, StoneDetector finds structural code clones based on its applied control flow representation of dominator trees, which are missed by CCAligner (21).
+Remaining code clones were classified as of other clone types (18) and false positives (9).
+In contrast, CCAligner’s code window approach allowed for finding subclones with many concentrated code modifications which are missed by StoneDetector (26).
+CCAligner, in addition, considers a function inside an anonymous class and the surrounding code as potential code clones, which is not the case for StoneDetector (31).
+The remaining code clones exclusively detected by CCAligner were classified as of other types (33) and false positives(10).
+StoneDetector vs Oreo
+Oreo uses functional metrics for clone detection, which determine, for example, the number of field accesses, array accesses, loops, etc., in code snippets.
+Clone detection in Oreo does not depend on the order of the instructions, but instead analyzes code properties by means of these functional metrics.
+In the manual review of the random samples, we found many structural clones (49) and subclones (14) that were overlooked by Oreo but detected by StoneDetector.
+These were often code clones that were very different in structure, which StoneDetector recognizes, but for which Oreo derives metric measures that are too distant from each other.
+The remaining clones found exclusively by StoneDetector were classified as of other clone types (31) and false positives (6).
+On the other hand, most of the code clones found exclusively by Oreo were structural clones (52) that often contained subpaths in internal substructures which were disordered or of very different lengths.
+It was also noticeable that Oreo thereof identified subclones (25) consisting of basic blocks, although they have very different lengths or their instructions were of different order, which is not a problem for Oreo due to its usage of functional metrics but pose an issues for the string metric comparison in StoneDetector.
+Remaining clones found exclusively by Oreo were assigned as of other types (21) or as false positives (2).
+StoneDetector vs All
+Finally, we also looked into the set of code clones in BigCloneBench which were exclusively identified by StoneDetector but not by the other clone detectors considered in our evaluation.
+From the total of 1,300,458 code clones identified by StoneDetector (cf.
+Table 4), a fraction of 235,982 clones were not detected by any of the other tools, i.e., almost each fifth code clone (18%).
+A closer examination of 100 randomly selected clones of this set again showed that StoneDetector finds significantly more subclones (49) and structural clones (35), which often differed greatly in size.
+Additional clones were classified as other (14) and false (2).
+4.3 Research Question 2: Generalizability
+Our second research question addresses the capability of clone detectors to identify code clones in source code with larger syntactical variability, in particular code clones of Strongly Type 3 and beyond.
+Besides, we are also interested in analyzing the clone detectors’ performance on benchmarks other than BigCloneBench in order to prevent the evaluation from being biased toward this particular benchmark, also considering the benchmark’s various issues as mentioned above.
+We therefore ran the clone detectors SourcererCC, NiCad, CCAligner, StoneDetector, NIL, CloneWorks, and Oreo on the three benchmarks GPTCloneBench, Google Code Jam, and Project CodeNet, respectively, and report the observed recall values in Table 5.
 We omitted Deckard and iClones due to their lower precision compared to the other tools.
 We again limit clones to those with a minimal size of 15 lines, use the same configurations, and depict recall in absolute and relative numbers.
-Benchmark SourcererCC NiCad CCAligner StoneDetector NIL Oreo
-GPTCloneBench 4,577 3,646 3,452 4,068 4,681 4,366
-26.9% 21.5% 20.3% 23.9% 27.6% 25.7%
-Google Code Jam 101,843 54,363 102,129 362,989 212,597 115,879
-0.9% 0.5% 0.9% 3.3% 1.9% 1.1%
-Project CodeNet 104,790 121,271 145,218 713,942 646,232 537,322
-0.9% 1.1% 1.4% 6.4% 5.8% 4.8%
+Benchmark SourcererCC NiCad CCAligner StoneDetector NIL Oreo CloneWorks
+GPTClone 4,577 3,646 3,452 4,068 4,681 4,366 3,987
+Bench 26.9% 21.5% 20.3% 23.9% 27.6% 25.7% 23.4%
+Google Code 101,843 54,363 102,129 362,989 212,597 115,879 91,794
+Jam 0.9% 0.5% 0.9% 3.3% 1.9% 1.1% 0.8%
+Project 104,790 121,271 145,218 713,942 646,232 537,322 266,888
+CodeNet 0.9% 1.1% 1.4% 6.4% 5.8% 4.8% 2.4%
 Table 5: Recall of clone detectors on GPTCloneBench, Google Code Jam, Project CodeNet
@@
 Thus, the recall values for all tools do not exceed 10% for both benchmarks.
-We can as well observe a distinction between the near-miss clone detectors NiCad and SourcererCC and tools addressing more difficult code clones, i.e., CCAligner, NIL, Oreo, and StoneDetector, where the latter ones perform better.
+We can as well observe a distinction between the near-miss clone detectors NiCad, CloneWorks, and SourcererCC and tools addressing more difficult code clones, i.e., CCAligner, NIL, Oreo, and StoneDetector, where the latter ones perform better.
 While the contrast is not so apparent on the benchmark GPTCloneBench, due to its inclusion of Type 1/2 clones, it becomes obvious for the other two benchmarks.
@@
 StoneDetector in particular excels at finding source code clones with larger syntactical variance when compared to other conventional state-of-the-art clone detectors on different benchmarks.
-Tool 1 million LOC 10 million LOC 100 million LOC
-NiCad 57s 18.49min 2.45h
-CCAligner 18s 6.58min -
-SourcererCC 4.17min 28.09min 46.52h
-iClones 24s 7.26min -
-StoneDetector 24s 6.17min 7.13h
-Oreo 2.42min 9.54min 7.10h
+Tool 1 million LOC 10 million LOC 100 million LOC 328 million LOC
+NiCad 57s 18.49min 2.45h -
+CCAligner 18s 6.58min - -
+SourcererCC 4.17min 28.09min 46.52h -
+iClones 24s 7.26min - -
+StoneDetector 24s 6.17min 7.13h 79h
+StoneDetector 24s 2.48min 1.11h 12.32h
+Oreo 2.42min 9.54min 7.10h 23.53h
+NIL 2s 17s 10.21min 1.44h
+CloneWorks 10s 1.46min 25.35min 4.07h
 Table 6: Scalability experiments on IJaDataset 2.0
-4.5 Research Question 4: Scalability
+4.4 Research Question 3: Scalability
 Scalability is another important requirement for clone detectors, i.e., such tools need to be able to process large quantities of code in reasonable time.
@@
 Sect.
-4.7).
-Due to these encouraging results, we performed another scalability check using StoneDetector on the entire IJaDataset 2.0 [45], comprising approx.
+4.5).
+Due to these encouraging results, we performed another scalability check using StoneDetector and the other clone detectors on the entire IJaDataset 2.0 [7], comprising approx.
 2.5 million Java source files with 328 million lines of code.
+The respective runtime results are as well shown in Fig.
+6.
+As can be observed, only four of the tools have been able to analyze the whole dataset.
+While CCAligner and iClones already did not scale for 100 million lines of code, NiCad failed for 328 million lines of code due to an internal threshold of input size and SourcererCC was cut off after running for a whole week.
+Overall, NIL and CloneWorks provide the best runtime, with NIL only requiring less than 2 hours for analyzing the whole IJaDataset.
 StoneDetector has been able to analyze this very large dataset in 79 hours.
-Again note that we report on an even faster version of StoneDetector when the Hamming distance is used as a string metric, as described in Sect.
-4.7.
-StoneDetector scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art.
-4.6 Research Question 5: Configuration
-SourcererCC NiCad CCAligner StoneDetector NIL Oreo
-Recall 35,843 35,847 35,744 35,843 35,783 35,480
-T 1 99.9% 99.9% 99.6% 99.9% 99.8% 98.9%
-Recall 4,432 4,550 4,530 4,579 4,421 4,506
-T 2 96.7% 99.3% 98.8% 99.9% 96.4% 98.3%
-Recall 3,814 4,098 4,055 4,140 3,892 3,893
-VST 3 91.6% 99.2% 97.4% 99.4% 93.5% 93.5%
-Recall 9,026 10,680 10,534 13,008 10,083 12,625
-ST 3 59.2% 70.0% 69.1% 85.3% 66.1% 82.8%
-Recall 3,935 476 8,099 14,537 8,548 22,965
-MT 3 4.8% 0.6% 10.0% 17.9% 10.5% 28.2%
-Recall 1,690 11 12,673 11,612 14,500 52,839
-T 4 0.0% 0.0% 0.2% 0.1% 0.2% 0.7%
+Due to its focus on finding harder-to-detect code clones, StoneDetector thus trades higher analysis time for recall.
+However, note that we report on a faster version of StoneDetector when the Hamming distance is used as a string metric with MD5 hashing, as described in Sect.
+4.5, which makes it a fast near-miss clone detector.
+We have included the respective runtime for this version of StoneDetector in Fig.
+6 as well.
+As can be seen, while not as fast as NIL or CloneWorks, StoneDetector then can analyze 328 million lines of code in 12 and a half hours.
+Like discussed in Sect.
+4.5, this version of StoneDetector though yields significantly poorer recall values.
+StoneDetector scales to large codebases comprising multiple 100 million lines of code.
+Using the Hamming distance instead of the modified LCS metric makes StoneDetector behave like a fast near-miss clone detector.
+4.5 Research Question 4: Tool Configuration
+SourcererCC NiCad CCAligner StoneDetector NIL CloneWorks Oreo
+Recall 35,843 35,847 35,744 35,843 35,783 35,844 35,480
+T 1 99.9% 99.9% 99.6% 99.9% 99.8% 99.9% 98.9%
+Recall 4,432 4,550 4,530 4,579 4,421 4,550 4,506
+T 2 96.7% 99.3% 98.8% 99.9% 96.4% 99.3% 98.3%
+Recall 3,814 4,098 4,055 4,140 3,892 4,094 3,893
+VST 3 91.6% 99.2% 97.4% 99.4% 93.5% 98.3% 93.5%
+Recall 9,026 10,680 10,534 13,008 10,083 14,165 12,625
+ST 3 59.2% 70.0% 69.1% 85.3% 66.1% 92.9% 82.8%
+Recall 3,935 476 8,099 14,537 8,548 2,782 22,965
+MT 3 4.8% 0.6% 10.0% 17.9% 10.5% 3.4% 28.2%
+Recall 1,690 11 12,673 11,612 14,500 46 52,839
+T 4 0.0% 0.0% 0.2% 0.1% 0.2% 0.0% 0.7%
 Table 7: Recall of clone detectors on Java source code on BigCloneBench with a minimal code clone size of 6 lines of code (T 1 - Type 1, T 2 - Type 2, VST 3 - Very-Strongly Type 3, ST 3 - Strongly Type 3, MT 3 - Moderately Type 3, T 4 - Type 4)
@@
 In order to achieve this objective, certain parameters of our clone detection method have to be chosen.
-Thus, our fifth research question addresses experiments on optimal configuration parameters in StoneDetector, in particular considering the threshold values for syntactical similarity of code clones and their minimum size.
+Thus, our fourth research question addresses experiments on optimal configuration parameters in StoneDetector, in particular considering the threshold values for syntactical similarity of code clones, their minimum size, the used similarity measure, i.e., string metric, and hashing algorithm for comparing and encoding code fragments, respectively.
 In the experiments presented so far, we have chosen 15 lines as the minimum size of code clone fragments.
@@
 As expected, the shown recall values are in general higher compared to those in Table 4.
-This can be well explained by Java methods with only 6 lines of source code, where half of the lines are just taken by the method’s annotations, signature, and curly braces.
+This can well be explained by Java methods with only 6 lines of source code, where half of the lines are just taken by the method’s annotations, signature, and curly braces.
 We, though, do not observe substantial changes in the tools’ comparisons, besides Oreo now outperforming StoneDetector in the case of Moderately Type 3 code clones.
@@
 6.
+Note that we provide the detailed results as supplementary information online [^18] https://stonedetector.fmi.uni-jena.de/.
 As can be seen, comparable recall values have been obtained, while precision values range from 96% to 100%, with increasing values for increasing minimum sizes of code clones.
@@
 also Sect.
-2.5).
+3.5).
 In the experiments conducted so far, this threshold value has been set to 70%, i.e., distance value $\tau=0.3$, which on the one-hand side is comparable to the threshold values as used in other clone detectors and on the other-hand side, in our experience, provides an optimal tradeoff of high recall and precision.
@@
 The abstraction of paths in StoneDetector, e.g., the uniform treatment of variables and constants as discussed in Sect.
-2, can lead to several identical paths in one description set.
+3, can lead to several identical paths in one description set.
 As an example, consider two paths that only differ from each other in their last node, and where the respective nodes each represent assignments with which the same/different variables are assigned to different/the same constants.
@@
 25 to 20 minutes on the whole BigCloneBench.
-Clone detection on source code with StoneDetector achieves the best results using a similarity threshold of 70% ($\tau=0.3$).
-Increasing the minimal size of clones in general increases the precision of clone detectors.
-4.7 Research Question 6: Metrics and Hashing
 LCS with 4- LCSmodified Leven- Needleman- Hamming Hamming LCSmodified
@@
 also Sect.
-2.3).
+3.3).
 Furthermore, since the path comparison is working on hashed fingerprints of the paths, the utilized hashing mechanism may as well influence StoneDetector’s performance.
-For our sixth research question, we therefore want to specifically analyze the role of the string metric and hashing mechanism.
-In our previously described experiments, we used a combination of a variant of the Longest Common Subsequence (LCS)-based metric ($\delta_{LCS_{modified}}$) in combination with Locality-Sensitive Hashing (LSH).
-This is in contrast to the authors’ previous work in [21], where we used the standard LCS-based metric in combination with a 4-byte prime number-based hashing mechanism.
-In a first experiment, we considered the role of using different string metrics on clone detection in StoneDetector.
+We therefore also want to specifically analyze the role of the string metric and hashing mechanism.
+In our previously described experiments, we used a combination of a variant of the longest common subsequence-based metric ($\delta_{LCS_{modified}}$) in combination with Locality-Sensitive Hashing (LSH).
+This is in contrast to the authors’ previous work in [8], where we used the standard LCS-based metric in combination with a simple 4-byte prime number-based hashing mechanism.
+In a third experiment, we considered the role of using different string metrics on clone detection in StoneDetector.
 To this end, we have chosen a subset of the string metrics implemented in StoneDetector, i.e., the LCS-based metric, modified LCS-based metric, Levenshtein distance, Needleman-Wunsch, Hamming distance, and modified Hamming distance.
-The latter metric depicts a variant of the Hamming distance in which different path lengths are not penalized, in the sense that the difference between different paths is not included in the distance measure as a penalty.
+The latter metric depicts a variant of the Hamming distance in which different path lengths are not penalized, in the sense that the difference in length between different paths is not included in the distance measure as a penalty.
 We refer the interested reader to Sect.
@@
 Table 2).
-The results of this first experiment are shown in Table 8.
+Precision is again measured by manual validation of a random sample.
+The results of this third experiment are shown in Table 8.
 Note that we did not observe any differences in the results when using either MD5 or a simple 4-byte hashing and are therefore denoting the results for the standard LCS-based metric with 4-byte hashing for the sake of comparison with our previous work [8].
@@
 While the use of the modified Hamming distance results in improved recall values, it features the worst precision.
-However, we observed significantly shorter runtimes of StoneDetector in the case of using either variant of the Hamming distance.
-Using the Hamming distance resulted in a speed-up of approximately 3.5 compared to the other metrics.
+However, we observed significantly shorter runtimes of StoneDetector in the case of using the Hamming distance, as reported in Sect.
+4.4 and shown in Table 6.
 As a rule of thumb to be taken from these measurements, we can thus state that the Hamming distance should be preferred over the other metrics when searching exact and near-miss clones, i.e., Type 1 and Type 2 code clones, due to its impact on shortening StoneDetector’s runtime.
-In a second experiment, the hashing algorithm’s influence on StoneDetector’s performance is evaluated.
+In a fourth experiment, the hashing algorithm’s influence on StoneDetector’s performance is evaluated.
 Since the modified LCS-based metric has so far achieved the best results, we consider only this string metric in this experiment and compare the tool’s recall and precision when using either the MD5 hashing algorithm or Locality-Sensitive Hashing (LSH).
@@
 We have therefore chosen the modified LCS-based metric and LSH as StoneDetector’s default configuration in our evaluation.
-Clone detection on source code with StoneDetector achieves the best results using the modified Longest Common Subsequence-based string metric in combination with Locality-Sensitive Hashing.
+Clone detection on source code with StoneDetector achieves the best results using a similarity threshold of 70% ($\tau=0.3$).
+Increasing the minimal size of clones in general increases the precision of clone detectors.
+StoneDetector performs best using the modified longest common subsequence (LCS)-based string metric in combination with locality-sensitive hashing (LSH).
 Using the Hamming distance instead, though, allows for faster clone detection.
-4.8 Research Question 7: Java Bytecode
+4.6 Research Question 5: Java Bytecode
 StoneDetector (stack-based) StoneDetector (register-based) iClones (stack-based)
@@
 In contrast, the Jimple format is a register-based three-address code, which is generated by Soot out of the stack-based Bytecode for program analysis and transformation.
-We consider both formats in our evaluation and are interested in the their different suitability for code clone detection.
+We consider both formats in our evaluation and are interested in their different suitability for code clone detection.
 Evaluating clone detectors for Java Bytecode is unfortunately not as straightforward as for Java source code, because the state-of-the-art benchmark BigCloneBench consists of only plain source files and is missing their respective dependencies [94].
@@
 A register-based representation of Java Bytecode is better suited for their detection compared to a stack-based representation.
+4.7 Summarization
+Our evaluation of StoneDetector in comparison with state-of-the-art conventional clone detection tools illustrates its strengths and its limitations.
+While near-miss clone detectors like NiCad, SourcererCC, and CloneWorks excel in identifying Type 1 and Type 2 code clones with high precision and fast runtimes (cf.
+Sect.
+4.2, Sect.
+4.4), StoneDetector achieves in its default configuration competitive performance results, though not providing for similar fast clone detection runtimes.
+StoneDetector however demonstrates its capabilites in finding harder-to-detect clones of Type 3 and Type 4 with acceptable precision in reasonable time, also when compared to advanced and/or Deep Learning tools such as NIL and Oreo, as exemplified on benchmarks Google Code Jam and Project CodeNet where StoneDetector achieved the best overall recall (cf.
+Sect.
+4.3).
+The manual qualitative examination of code clones revealed that the tool in fact finds certain types which are not detected by any of the other tools (cf.
+Sect 4.2.1).
+Furthermore, StoneDetector can be adapted to other configurations (cf.
+Sect.
+4.5), e.g., using the Hamming distance metric for scalable and fast near-miss clone detection, which makes it a versatile and rich platform for Java code clone detection and experimentation, including Bytecode-based clone detection (cf.
+Sect.
+4.6).
+On the one hand, StoneDetector in its default configuration is thus most beneficial in software maintenance scenarios, where developers are challenged by growing inconsistencies across clone instances [118].
+Note that such syntactically more deviating clone instances have also been shown to be prone to delays in propagating code changes, referred to as late propagation [67], potentially resulting in bugs, which makes their detection and tracking even more important.
+In these cases, StoneDetector’s tolerance for refactored statements, i.e., structural code clones, and larger code edits or subsumed/embedded code fragments, i.e., subclones, enable it to capture clones and therefore bugs that other clone detectors may miss, as revealed in our evaluation.
+Such scenarios are especially relevant in software maintenance, where inconsistent and insufficient identification of code clones has been shown to correlate with fault proneness [38].
+StoneDetector’s ability to detect structurally modified code clones also supports consolidation of functional and architectural adaptations [9], which is less feasible with clone detectors optimized for exact or near-miss clones.
+On the other hand, StoneDetector allows for other configurations.
+For example, in certain scenarios, the primary objective of clone detection may be the rapid identification of exact duplicates or near-miss clones in large codebases.
+Note that such code clones often constitute a large fraction in practice [64].
+As mentioned above, using the Hamming distance metric and MD5 hashing makes StoneDetector behave like a fast near-miss clone detector, yielding a speed-up of approx.
+6.3, though trading accuracy for speed.
+Apart from StoneDetectur thus being adaptable to software maintenance scenarios, the tool provides a rich platform for Java code clone experimentation including various configuration parameters like, e.g., used distance metrics, hashing algorithms, code representations.
 5 Related Work
@@
 Following a recent survey on this research topic by Nasrabadi et al.
-[6], we differentiate text-, token-, tree-, graph-, and metric-based methods as well as hybrid, Deep Learning, and dynamic approaches.
+[70], we differentiate text-, token-, tree-, graph-, and metric-based methods as well as hybrid, Machine/Deep Learning, and dynamic approaches.
 For an extensive discussion of code clone techniques and tools, we refer the interested reader to one of the various surveys available on the topic [83, 70, 81, 2] and here provide a condensed overview of the most formative and relevant work.
@@
 Similarly, Johnson proposed to use the Karp-Rabin algorithm and sliding windows for hashed code lines, called fingerprints, to find duplicated code fragments [42].
-In order to also detect Type 2 clones, he removes whitespace and comments and maps all identifiers, keywords, and literals to a common placeholder.
+In order to also detect Type 2 clones, he also removes whitespace and comments and maps all identifiers, keywords, and literals to a common placeholder.
 The text-based clone detector Vuddy uses whole-function hashes on pre-processed source code for generating function fingerprints.
@@
 The code pre-processing can be adjusted to different programming languages using a customizable grammar definition.
-NiCad is using the Longest Common Subsequence algorithm for finding matching code and is one of the clone detectors used in our comparative evaluation in Sect.
+NiCad is using the longest common subsequence algorithm for finding matching code and is also one of the clone detectors used in our comparative evaluation in Sect.
 4.
@@
 Furthermore, modern token-based clone detectors typically scale up to large code bases with several hundred million lines of code [88].
-CCFinder [57] lexes two source code fragments into token sequences, thereby transforming some tokens according to certain rules, e.g., replacing mapping identifiers to a common token.
+CCFinder [44] lexes two source code fragments into token sequences, thereby transforming some tokens according to some rules, e.g., replacing mapping identifiers to a common token.
 The thus normalized token sequences are then handed over to the suffix tree algorithm for finding matching code fragments.
@@
 Scaling clone detectors for large codebases has also been the focus of subsequent research [57].
+CloneWorks also uses a token-based approach combined with several techniques to scale and fasten clone detection, including Jaccard similarity metric, sub-block filtering heuristics, index creation and partitioning [104].
 Another token-based technique is described by Ragkhitwetsagul and Krinke [80], their tool Siamese uses token sequences with different levels of abstractions and methods from document retrieval on their n-gram vector representations.
@@
 These so-called large-gap clone detectors often apply more sophisticated string metrics for aligning and matching token sequences, e.g., e-edit distance [110] or Smith-Waterman sequence alignment [63].
-We include the tools SourcererCC, iClones, and CCAligner as representative token-based clone detectors in our evaluation in Sect.
+We include the tools SourcererCC, iClones, CloneWorks, NIL, and CCAligner as representative token-based clone detectors in our evaluation in Sect.
 4
@@
 With the advent of multi-stage code clone detectors, metric-based techniques have also been proposed for pre-filtering candidate pairs of code fragments according to their metrics [87, 41].
-5.5 Deep Learning Techniques
+5.5 Machine/Deep Learning Techniques
 In the last decade, the application of machine learning and Deep Learning has become more widespread in clone detection.
@@
 The ability to thus automatically learn and recognize the hidden patterns of code clones from, e.g., today’s large and omnipresent code repositories, should guarantee promising results.
-While in particular Deep Learning clone detectors have shown very good performance in the detection of Type 1 to Type 4 code clones lately, research has recently also highlighted common problems and pitfalls in the application of Deep Learning to clone detection, i.e., training/evaluation bias, overfitting, and missing ground truth [78, 79, 80].
+While in particular Deep Learning clone detectors have shown very good performance in the detection of Type 1 to Type 4 code clones lately, research has recently also highlighted common problems and pitfalls in the application of Deep Learning to clone detection, i.e., training/evaluation bias, overfitting, and missing ground truth [53, 92, 52, 21].
 Prior approaches considered conventional machine learning methods.
@@
 Similarly, Zhang et al.
-are using, in their architecture ASTNN, bidirectional recurrent units on embedded statement trees for encoding abstract syntax trees and identifying code clones [86].
+are using in their architecture ASTNN bidirectional recurrent units on embedded statement trees for encoding abstract syntax trees and identifying code clones [121].
 Different settings for embeddings of abstract syntax trees are evaluated by Büch and Andrzejak [13], where they use a siamese architecture that optimizes the embeddings of two abstract syntax trees such that the cosine similarity in the case of code clones is maximized.
@@
 The resulting encodings of code fragments can then be compared via their Euclidean distances.
+Hu et al.
+[37] choose manual feature engineering to encode abstract syntax trees using different centrality values and then employ decision trees for finding matching code clones based on the resulting vector representations.
+Different similarity measures based upon code fragments’ token sequences are considered by Feng et al.
+[28], which feed the resulting feature vectors into several classical machine learning algorithms for detecting code clones.
 The Siamese architecture is also used in the hybrid clone detector Oreo, where the model is used with metrics-based feature vectors for detecting harder-to-find clones, achieving in particular good results for Moderately Type 3 clones and beyond [87].
@@
 Dynamic methods focus on the computations of code instead of its syntax and usually target binary code.
-Therefore, dynamic analyses or testing is used to execute code fragments with sample inputs and to track their behavior.
+Therefore, dynamic analyses or testing is used to execute code fragments with sample inputs and track their behavior.
 The resulting data can then be used to match semantically similar code and, in this way, in particular, to detect Type 4 code clones.
@@
 In this way, our method is different from most other conventional approaches to clone detection for Java, since the approach incorporates control flow besides syntax and, as a result, achieves better performance when asked to find code clones with larger syntactical variance, i.e., Type 3 and Type 4 clones.
-The method for clone detection has been implemented in the publicly available StoneDetector platform, which provides a versatile system for finding code clones in both Java source code and Bytecode, as well as supporting manifold experiments on clone detection’s parameters.
+The method for clone detection has been implemented in the publicly available StoneDetector tool, which provides a versatile system for finding code clones in both Java source code and Bytecode, as well as supporting manifold experiments on clone detection’s parameters.
 Using the implementation in StoneDetector, we conducted a thorough evaluation of our approach with respect to the state-of-the-art benchmarks for Java code clone detection, i.e., BigCloneBench, Google Code Jam, GPTCloneBench, and Project CodeNet.
-Furthermore, we experimentally compared the performance of the code clone detection approach in StoneDetector to other state-of-the-art clone detectors, including NiCad, iClones, CCAligner, SourcererCC, Deckard, Oreo, and NIL.
+Furthermore, we experimentally compared the performance of the code clone detection approach in StoneDetector to other state-of-the-art clone detectors, including NiCad, iClones, CCAligner, SourcererCC, Deckard, Oreo, CloneWorks and NIL.
 The experiments proved StoneDetector’s competitive capabilities in finding exact near-miss clones with respect to precision and recall.
@@
 Further experiments demonstrated StoneDetector’s ability to analyze large code bases with up to multiple million lines of code.
-Besides its performance, we highlight the versatility of the platform for experimenting with different parameters for code clone detection, including the similarity threshold, minimal size of code clones, string metrics, e.g., Hamming distance, Longest Common Subsequence, and hashing methods, e.g., Locality-Sensitive Hashing.
+Besides its performance, we highlight the versatility of StoneDetector experimenting with different parameters for code clone detection, including the similarity threshold, minimal size of code clones, string metrics, e.g., Hamming distance, Longest Common Subsequence, and hashing methods, e.g., Locality-Sensitive Hashing.
 Eventually, StoneDetector’s ready-made extensibility to other programming languages and program representations is presented, as is shown for stack-based and register-based Bytecode formats.
-While this paper focuses on conventional code clone detection and therefore omits an in-depth discussion and comparative evaluation with Deep Learning approaches to code clone detection, we want to analyze and incorporate such models into the StoneDetector platform in our future work.
+While this paper focuses on conventional code clone detection and therefore omits an in-depth discussion and comparative evaluation with Deep Learning approaches to code clone detection, we want to analyze and incorporate such models into StoneDetector in our future work.
 While we acknowledge the promising results of, in particular, large language models in recent work [25], as discussed in the related work section, we also assume that conventional methods of clone detection conjoined with their advantages, as presented in this paper, will play a role in the future.
 We though concede that the assessment of the advantages and disadvantages of conventional code clone detection in comparison with Deep Learning approaches, as well as their incorporation into a possible combined approach, requires further research and analysis.
-Besides, we are also interested in extending the StoneDetector platform with respect to other programming languages, which have not been in the research focus for code clone detection.
+Besides, we are also interested in extending StoneDetector with respect to other programming languages, which have not been in the research focus for code clone detection.
+Finally, we believe that StoneDetector’s runtime can be improved by further optimizing its implementation as well as examining configurations which trade off accuracy against better runtime, as has been discussed for the Hamming distance.